### v1

# TrinityDNA: A Bio-Inspired Foundational Model for Efficient Long-Sequence DNA Modeling

Abstract
The modeling of genomic sequences presents unique challenges due to their long length and structural complexity. Traditional sequence models struggle to capture long-range dependencies and biological features inherent in DNA. In this work, we propose TrinityDNA, a novel DNA foundational model designed to address these challenges. The model integrates biologically informed components, including Groove Fusion for capturing DNA’s structural features and Gated Reverse Complement (GRC) to handle the inherent symmetry of DNA sequences. Additionally, we introduce a multi-scale attention mechanism that allows the model to attend to varying levels of sequence dependencies, and an evolutionary training strategy that progressively adapts the model to both prokaryotic and eukaryotic genomes. TrinityDNA provides a more accurate and efficient approach to genomic sequence modeling, offering significant improvements in gene function prediction, regulatory mechanism discovery, and other genomics applications. Our model bridges the gap between machine learning techniques and biological insights, paving the way for more effective analysis of genomic data. Additionally, we introduced a new DNA long-sequence CDS annotation benchmark to make evaluations more comprehensive and oriented toward practical applications.
1 Introduction
The rapid advancements in large-scale, long-sequence modeling, particularly in the realm of Natural Language Processing (NLP) (team2023gemini, ; achiam2023gpt, ), have radically transformed the way we approach complex data. Deep learning models, such as Transformers (vaswani2017attention, ), have achieved unprecedented success in tasks that span from language translation to text generation, revolutionizing not just NLP but a variety of other fields. These models have proven their capability to capture intricate dependencies in data, providing solutions to challenges that were once considered insurmountable. With these breakthroughs in NLP, there has emerged an exciting opportunity to extend the power of sequence modeling to a completely different domain—genomics—where data shares some key similarities, such as its sequential nature (zhou2015predicting, ).
Genomic data, particularly DNA sequences, consists of extraordinarily long strings of information that encode the fundamental building blocks of life. Unlike the highly dense and structured data typically encountered in NLP, genomic sequences are sparse in nature, containing vast stretches of repetition and variability (liu2024genbench, ). Despite this, they hold a rich repository of biological information that is crucial for understanding gene functions, regulatory mechanisms, and cellular processes. The ability to model DNA sequences deeply could lead to breakthrough applications in personalized medicine, genetic engineering, and the overall understanding of biological systems. However, effectively capturing the dependencies within such long, sparse sequences remains a significant challenge.
Figure 1: Overview of TrinityDNA Model: (Left) The evolutionary training strategy of TrinityDNA, progressing from prokaryotic DNA to multi-species eukaryotic DNA, and its DNA-targeted long-sequence modeling approach addressing structural features such as bidirectional complementarity and major/minor grooves. (Right) Radar chart illustrating the state-of-the-art performance on the zero-shot performance of our models versus popular models such as EVO and Caduceus. The TrinityMicroDNA model refers to a model trained only on prokaryotic data, which typically performs better on prokaryotic tasks, while TrinityDNA is better on eukaryotic tasks.
While the parallels between NLP and genomics are evident, directly applying traditional NLP models to genomic sequences proves difficult. The sparse, low-density nature of DNA sequences means that existing models often struggle to identify long-range dependencies and interpret the underlying biological structures(mallet2021reverse, ; zhou2021towards, ). Moreover, the lack of biologically informed features in current models limits their effectiveness in genomic contexts. Many models trained on single-species data perform poorly when generalized to other species or broader biological contexts. As a result, the impact of these models on genomic research has been somewhat restricted, and their applicability to real-world challenges remains limited.
To address these issues, we introduce TrinityDNA, a novel DNA foundational model specifically designed to overcome the current limitations of genomic sequence modeling, as shown in Fig 1. TrinityDNA leverages the latest advancements in deep learning to create a model that is optimized for the unique challenges posed by DNA sequences while also incorporating key biological insights. The contributions are listed as follows:
  • Bio-inspired Design: A multi-level architecture that is optimized for DNA sequences and leverages the Groove Fusion module and Reverse Complement (RC) fusion strategy. This design captures and exploits the unique structural properties of DNA, enabling long bi-directional genomic modeling.
  • Evolutionary Training Strategy: A multi-species training regimen that spans a variety of organisms from prokaryotes to eukaryotes, enabling the model to generalize different genomic contexts and sequence lengths.
  • Comprehensive Large-Scale Data Integration: Curated and integrated datasets from prominent genomic databases such as GTDB, IMG, and RefSeq, ensuring a diverse and high-quality foundation for model training.
  • New Benchmark for Long-Sequence Inference: We introduce a novel CDS Annotation Benchmark that focuses on gene-structure labeling in prokaryotic genomes, assessing both long-sequence modeling and practical annotation performance.
In summary, TrinityDNA represents a significant advancement in DNA sequence modeling, addressing critical limitations of existing models through innovative architectural designs, incorporation of biological insights, and robust training strategies. By bridging the gap between NLP advancements and genomic data intricacies, TrinityDNA paves the way for more accurate predictions of gene functions, enhanced understanding of genetic regulation mechanisms, and broader applications in genomics research.
2 Background
2.1 DNA Terminology for Structures.
Basic Composition
Deoxyribonucleic acid (DNA) is the hereditary material in most living organisms, consisting of long sequences of nucleotides. Each nucleotide comprises a phosphate group, a deoxyribose sugar, and one of four nitrogenous bases: adenine (A), thymine (T), cytosine (C), and guanine (G). A DNA sequence can be represented as a string $S=(s_{1},s_{2},\ldots,s_{N})$, where $s_{i}\in\{A,T,C,G\}$ and $N$ denotes the sequence length. Genomic sequences are characterized by their low-density encoding and considerable length, often extending to millions of nucleotides, which presents unique challenges for sequence modeling.
Minor and Major Grooves.
The double helix structure of DNA features two distinct grooves: the minor groove and the major groove. These grooves arise from the asymmetric positioning of the phosphate backbone relative to the base pairs. (1) Major Groove: wider and deeper, the major groove provides greater accessibility for protein binding and molecular interactions. It generally covers five to seven nucleotides. (2) Minor Groove: narrower and shallower, the minor groove presents a different arrangement of hydrogen bond donors and acceptors. While less accessible than the major groove, its length is three to five nucleotides. Understanding the geometric properties of these grooves is essential for predicting functional regions within the genome.
Reverse Complement Strands.
DNA molecules consist of two complementary strands running in opposite directions, a feature known as antiparallel orientation. For a DNA sequence $S=(s_{1},s_{2},\ldots,s_{N})$, its reverse complement $S^{R}$ is defined as:
$S^{R}=(s_{N}^{C},s_{N-1}^{C},\ldots,s_{1}^{C})$
where $s_{i}^{C}$ denotes the complementary base of $s_{i}$ following the base-pairing rules: $A\leftrightarrow T$ and $C\leftrightarrow G$. This reverse complementarity is fundamental to DNA replication and transcription processes. Incorporating reverse complement information into computational models enhances their ability to capture symmetrical and complementary patterns, thereby improving predictions related to gene annotation and regulatory element identification.
2.2 DNA Long-Sequence Modeling
Structured State Space Models (SSMs).
A prominent class of models for handling long-range dependencies is based on Structured State Space Models (SSMs) (gu2021efficiently, ; gu2021combining, ; gu2022parameterization, ; gupta2022diagonal, ; smith2022simplified, ; dao2022hungry, ). These models emerge from discretizing a continuous-time linear system:
${{\bm{h}}}(t)$ $={\bm{A}}{\bm{h}}(t)+{\bm{B}}\,x(t),\quad$ $y(t)={\bm{C}}\,{\bm{h}}(t)+{\bm{D}}\,x(t),$
${\bm{h}}_{t+1}$ $=\overline{{\bm{A}}}\,{\bm{h}}_{t}+\overline{{\bm{B}}}\,x_{t},\quad$ $y_{t+1}={\bm{C}}\,{\bm{h}}_{t}+{\bm{D}}\,x_{t},$
where ${\bm{h}}(t)$ (or ${\bm{h}}_{t}$) is an internal state that can capture long-range dependencies in the input sequence $x(t)$ (or $x_{t}$). Through efficient convolution-based implementations of these recurrences, SSMs have shown strong performance on very long sequences. However, conventional SSMs do not explicitly adapt their parameters to specific tokens or positions, which may limit expressivity for tasks such as DNA sequence modeling. Therefore, a line of SSMs are designed for long-sequence DNA.
HyenaDNA (EVO).
HyenaDNA (poli2023hyena, ) is an SSM variant, a decoder-style model capable of handling long genomic sequences (e.g., hundreds of thousands of tokens). It leverages the Hyena operator, which replaces traditional attention with a fast convolution-based mechanism. Concretely, HyenaDNA blocks compute a Toeplitz convolution filter on projected input tokens, allowing processing of very large contexts (in $\mathcal{O}(L\log L)$ time) without the quadratic cost typically associated with attention.
Caduceus (MambaDNA).
MambaDNA (schiff2024caduceus, )—also referred to as Caduceus—builds on the selective SSM approach of Mamba and incorporates reverse-complement symmetry, a core property of DNA sequences. In standard Mamba, the module processes sequences in a single direction. MambaDNA extends this design in two ways: (1) BiMamba: Instead of only left-to-right processing, BiMamba applies the Mamba block twice. (2) RC Equivariance: MambaDNA explicitly enforces RC symmetry by taking a sequence and its RC as inputs to the same SSM-based module. See more details in Appendix D.
3 TrinityDNA: Sequence, Structure, and Strategy for DNA Modeling
[…]
3.1 Preliminaries
Lost in the locality.
While SSMs theoretically excel at handling long sequences, they inherently exhibit a locality bias (wang2024understand, ). This issue is amplified in DNA sequence modeling, where dependencies across vast stretches of genetic material need to be captured for accurate biological interpretation. Specifically, in genomic data, long-range dependencies often span tens of thousands or even hundreds of thousands of base pairs. Existing models like SSMs typically focus on local dependencies due to the computational challenges of processing long sequences. Consequently, our empirical results in Figure 2 demonstrate that the SSM-based model (Caduceus) lost their focus as sequence length increased, while the full-attention-based model (DNABERT2) suffered from heavy computation. These limitations hamper the ability of modern DNA foundation models to fully capture the complexities inherent in DNA sequences, particularly when dealing with biological functions and regulatory mechanisms that depend on large, interconnected genomic regions.
Figure 3: Average attention entropy of full self-attention models as sequence length increases.
Oversmoothing in long attention.
As sequences grow in training, full self-attention “flattens out”: attention scores converge toward a uniform, high-entropy distribution in Figure 3, so every token is weighted almost equally, and useful signals vanish. This oversmoothing hits low-density data—images, DNA-hardest, where informative tokens are sparse and far apart. In shorter windows, the same model still shows specialized heads, e.g., retrieval vs. induction (xiao2024duo, ), but that diversity collapses at the kilobase scale. These findings motivate a multi-window, multi-head scheme that mixes narrow local windows with broader global ones to curb entropy and retain head specialization.
Figure 4: Model Architecture of TrinityDNA: The model integrates DNA sequences and structural features by considering its grooves and reverse complementary sequence with shared parameters.
3.2 Groove Fusion Module
To account for the minor and major grooves in DNA sequences, we propose a Groove Fusion module that combines convolutional operations of varying kernel sizes. These grooves have distinct structural and functional roles in DNA, with the major groove offering greater accessibility for protein binding and the minor groove being involved in different molecular interactions. To model these differences, we perform tokenization on the DNA sequence using three convolutional kernels of sizes 3, 5, and 7. This multi-scale convolution approach enables the model to focus on different spatial features across the sequence, effectively capturing the structural nuances associated with the two grooves.
Formally, the Groove Fusion process can be defined as:
$\text{GrooveFusion}(S)=\sum_{k\in\{3,5,7\}}\text{GELU}(\text{Conv}_{k}(S))$
where $\text{Conv}_{k}$ represents the convolution operation with kernel size $k$, and $S$ is the input DNA sequence. The output of each convolution operation is fused to capture the multi-scale contextual information necessary for interpreting the structural variations between the major and minor grooves.
3.3 Sliding Multi-Window Attention
To overcome the locality bias and attention oversmoothing, we revisit the design of Multi-Head Attention (MHA), focusing on varying attention window sizes for different heads. In traditional MHA, each head typically attends to the entire sequence using a fixed attention window, which can limit the model’s ability to capture dependencies at different scales. In the context of DNA sequence modeling, dependencies are not uniform across the sequence. For example, short motifs require local attention, while regulatory regions may span longer stretches of DNA. To address this, we introduce a multi-scale attention grouping strategy named SMWA. In this approach, we assign different attention heads to focus on different scales of dependencies in the DNA sequence, allowing the model to specialize in either local or global dependencies depending on the scale of the feature being modeled. Formally, we define the window sizes for each attention head $h$ as $L_{h}\in\mathbb{N}$, with $L_{h}$ representing the length of the attention window for head $h$. For head $h$, the attention mechanism is computed using a sliding window of size $L_{h}$ over the sequence, allowing the model to focus on different scales. Specifically, the attention calculation is constrained to the sliding window around each position, which captures dependencies within a local region of the sequence. The sliding window can be defined as:
$\text{Attn}_{h}(S_{i})=\text{Softmax}\left(\frac{Q_{h}(i)K_{h}(i+[-L_{h},L_{h}])^{T}}{\sqrt{d_{k}}}\right)V_{h}(i+[-L_{h},L_{h}])$
where $Q_{h}\in\mathbb{R}^{N\times d_{k}}$ is the query matrix for head $h$, $K_{h}\in\mathbb{R}^{N\times d_{k}}$ is the key matrix, $V_{h}\in\mathbb{R}^{N\times d_{v}}$ is the value matrix, $N$ is the sequence length, and $d_{k}$ and $d_{v}$ are the dimensionality of the key and value vectors, respectively. $i$ is the index of the sequence, and $[-L_{h},L_{h}]$ represents the range of indices for the sliding window around $i$. This enables each head to specialize in attending to either short-range or long-range dependencies by adjusting $L_{h}$. The final output of the multi-head attention layer is the concatenation of all heads’ outputs, followed by a linear transformation:
$\text{SMWA}(S)=\text{Concat}(\text{Attn}_{1},\text{Attn}_{2},\dots,\text{Attn}_{H})W_{O}$
where $W_{O}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ is a learned output weight matrix, $H$ is the number of attention heads, and $d_{\text{out}}$ is the output dimensionality. This multi-scale attention mechanism allows the model to simultaneously capture both local and global dependencies by allocating different heads to focus on different sequence scales. Thus, shorter sequences can be captured by heads with smaller windows, while longer-range dependencies can be modeled by heads with larger windows, enabling the model to capture the hierarchical nature of DNA sequences better.
3.4 Gated Reverse Complement
To leverage the reverse complementarity inherent in DNA sequences, we introduce a novel Gated Reverse Complement (GRC) mechanism. This mechanism employs a shared Transformer module to process both the forward and reverse complement sequences in parallel. The reverse complement of a DNA sequence $S=(s_{1},s_{2},\dots,s_{N})$ is defined as $S^{R}=(s_{N}^{C},s_{N-1}^{C},\dots,s_{1}^{C})$, where $s_{i}^{C}$ denotes the complementary base of $s_{i}$ (with base-pairing rules $A\leftrightarrow T$, $C\leftrightarrow G$). The GRC mechanism works by feeding both the forward and reverse complement sequences into a shared SMWA-equipped Transformer network $f_{\theta}$, where the outputs are gated using a linear gating mechanism to combine the two representations effectively. The gating at the final layer is given by:
$\text{Output}=\text{GRC}(S,S^{R})=$ $f_{\theta}(S)+\sigma(W_{G}\cdot f_{\theta}(\text{Flip}(S^{R})))$
where $W_{G}$ are learned weights, $\sigma$ represents the sigmoid or identity function, and the Flip operator means to reverse the sequence as the original order. This allows the model to simultaneously learn the representations of both forward and reverse sequences, capitalizing on the symmetry of DNA, which is crucial for many biological processes.
3.5 Evolutionary Training Strategy
The Evolutionary Training Strategy (ETS) approach leverages a two-stage, evolution-inspired training strategy to progressively address the varying complexities of genomic data. In the first stage, the model is trained on prokaryotic genomes, which average around 924 base pairs (xu2006genome, ; nguyen2024sequence, ) and present relatively straightforward regulatory architectures. Through this foundational phase, the model captures essential DNA sequence motifs and organizational patterns. Subsequently, the second stage introduces eukaryotic genomes, known for their intron-exon structures and gene lengths that can span tens of kilobases (dalla2024nucleotide, ). Alongside this transition, the model’s context window is enlarged from 8K to 100K base pairs, accommodating multiple co-expressed genes and complex regulatory elements.
Figure 5: Scaling Behaviors of Our Proposed Model. (Left) Evaluation perplexity (PPL) against total FLOPs across multiple architectures, showing consistent improvements to various baselines. (Right) Impact of increasing context length (8k, 30k, 100k) on a eukaryotic dataset, where PPL steadily decreases with longer context windows.
4 Experiments
4.1 Pre-training
Data.
We adopt masked language modeling (MLM) (devlin2018bert, ) and character-level tokenization for all our DNA models. Following our Evolutionary training strategy, we conduct pre-training in two phases: Stage 1 (Short-Sequence Pre-training): We use prokaryotic (bacterial and archaeal) DNA data from the OpenGenome dataset introduced in nguyen2024sequence , training on sequences of length 8k to learn fundamental nucleic acid patterns in shorter contexts. Stage 2 (Multi-species Post-training): We then continue pre-training on a multi-species collection, the Multispecies dataset, presented in dalla2024nucleotide . The sequences in this stage can be as long as 100k, spanning archaebacteria, fungi, vertebrate genomes, and more. This step exposes the model to a rich spectrum of evolutionary signals, enabling it to handle much longer contexts and to capture the diverse structural intricacies of eukaryotic DNA. Hence, we propose two main models, TrinityMicroDNA and TrinityDNA, each with 1 billion (1B) parameters. The former is trained solely on prokaryotic data, while the latter builds upon the former by post-training on eukaryotic data. A “100k” label indicates an increased context window size of up to 100k. By transitioning from bacterial genomes to longer eukaryotic genomes, our method achieves broad coverage of genomic features across scales. See more details in Appendix A.
Scaling Laws.
Figure 5 illustrates three key aspects of our model’s scaling behavior across parameter sizes, pre-training context lengths (8k, 30k, 100k). (Left) We plot the compute-perplexity frontier against total FLOPs for several architectures, demonstrating that our approach consistently outperforms baseline methods (Transformer, Caduceus, EVO, and EVO2) at every compute level in different parameter sizes (6M to 1B). (Right) We examine the effect of increasing context window sizes on a eukaryotic benchmark, finding a steady drop in perplexity when moving from 8k to 30k, and a further substantial improvement at 100k.
Figure 6: Comprehensive ablation study and efficiency analysis on TrinityDNA.
4.2 Ablation Study and Analysis
Table 1: Effect of GRC, GFM, and SMWA on pre-training perplexity.
Components W/O W
GRC 2.731 2.599 (-0.132)
GFM 2.599 2.534 (-0.065)
SMWA 2.534 2.544 (+0.010)
We comprehensively analyze our model by answering three key questions on the 470M scale: (1) Are the proposed modules effective? (2) How efficient of TrinityDNA for long-sequence DNA computing? (3) Why do we use ETS for DNA modeling?
(1) Effectiveness.
Table 1 shows that GFM consistently lowers perplexity by modeling the spatial ‘groove’ features of DNA sequences. Likewise, incorporating GRC (which captures reverse-complement patterns) yields a notable drop in PPL, reflecting the importance of complementary-strand information. Meanwhile, SMWA enables multi-scale context handling, trading off some computational overhead for competitive perplexity across local and longer-range dependencies. SMWA also brings some performance gain due to the low-density perplexity of the nucleotide sequence.
(2) Efficiency.
The left panel of Figure 6 contrasts token-throughput as we sweep both sequence length and micro-batch size. Across every setting, TrinityDNA sits clearly at the top, retaining more than 80% of its short-sequence throughput even at 64 k tokens. This robustness comes from its sliding multi-window attention and optimized fused kernels, which keep memory traffic almost flat as context grows. By contrast, DNABERT-2 exhibits the steepest decline: its throughput falls by nearly an order of magnitude once the sequence length exceeds 16 k, reflecting the quadratic cost of its vanilla soft-max attention. The remaining baselines lie between these two extremes, showing gradual slowdowns that track their respective attention or RNN bottlenecks.
(3) Training Strategies.
The right plot in Figure 6 shows a comprehensive ablation study separating the contributions of dataset size and evolutionary training strategy. The table shows perplexity scores across different model configurations. Key finding: Models initialized with weights from prokaryotic pre-training and then fine-tuned on combined data show better performance than models trained from scratch on the combined dataset. This validates both the importance of large, diverse datasets and our evolutionary training approach.
4.3 Downstream Tasks
We employ TrinityMicroDNA and TrinityDNA with 1B parameters as our base model for downstream evaluation, and we use LoRA tuning except for the zero-shot task.
4.3.1 GUE Benchmark
Table 2: GUE Benchmark performance comparison across various datasets and tasks.
Tasks DNABERT NT DNABERT2 Caduceus HyenaDNA TrinityDNA
# Params 86M 2.5B 117M 40M 1.6M 1B
H3 0.731 ±0.015$\pm$ 0.015± 0.015 0.788 ±0.012$\pm$ 0.012± 0.012 0.783 ±0.014$\pm$ 0.014± 0.014 0.799 ±0.029$\pm$ 0.029± 0.029 0.779 ±0.037$\pm$ 0.037± 0.037 0.814 ±0.014$\pm$ 0.014± 0.014
H3K14ac 0.401 ±0.018$\pm$ 0.018± 0.018 0.562 ±0.015$\pm$ 0.015± 0.015 0.526 ±0.019$\pm$ 0.019± 0.019 0.541 ±0.212$\pm$ 0.212± 0.212 0.612 ±0.065$\pm$ 0.065± 0.065 0.694 ±0.016$\pm$ 0.016± 0.016
H3K36me3 0.473 ±0.017$\pm$ 0.017± 0.017 0.620 ±0.012$\pm$ 0.012± 0.012 0.569 ±0.015$\pm$ 0.015± 0.015 0.609 ±0.109$\pm$ 0.109± 0.109 0.613 ±0.041$\pm$ 0.041± 0.041 0.692 ±0.014$\pm$ 0.014± 0.014
H3K4me1 0.414 ±0.012$\pm$ 0.012± 0.012 0.553 ±0.011$\pm$ 0.011± 0.011 0.505 ±0.015$\pm$ 0.015± 0.015 0.488 ±0.102$\pm$ 0.102± 0.102 0.512 ±0.024$\pm$ 0.024± 0.024 0.611 ±0.015$\pm$ 0.015± 0.015
H3K4me2 0.323 ±0.014$\pm$ 0.014± 0.014 0.365 ±0.010$\pm$ 0.010± 0.010 0.311 ±0.013$\pm$ 0.013± 0.013 0.388 ±0.101$\pm$ 0.101± 0.101 0.455 ±0.095$\pm$ 0.095± 0.095 0.480 ±0.013$\pm$ 0.013± 0.013
H3K4me3 0.278 ±0.015$\pm$ 0.015± 0.015 0.403 ±0.011$\pm$ 0.011± 0.011 0.363 ±0.014$\pm$ 0.014± 0.014 0.440 ±0.202$\pm$ 0.202± 0.202 0.549 ±0.056$\pm$ 0.056± 0.056 0.522 ±0.015$\pm$ 0.015± 0.015
H3K79me3 0.612 ±0.013$\pm$ 0.013± 0.013 0.647 ±0.010$\pm$ 0.010± 0.010 0.674 ±0.012$\pm$ 0.012± 0.012 0.676 ±0.026$\pm$ 0.026± 0.026 0.672 ±0.048$\pm$ 0.048± 0.048 0.741 ±0.014$\pm$ 0.014± 0.014
H3K9ac 0.512 ±0.015$\pm$ 0.015± 0.015 0.560 ±0.013$\pm$ 0.013± 0.013 0.556 ±0.017$\pm$ 0.017± 0.017 0.604 ±0.048$\pm$ 0.048± 0.048 0.581 ±0.061$\pm$ 0.061± 0.061 0.662 ±0.014$\pm$ 0.014± 0.014
H4 0.793 ±0.012$\pm$ 0.012± 0.012 0.817 ±0.015$\pm$ 0.015± 0.015 0.807 ±0.011$\pm$ 0.011± 0.011 0.789 ±0.020$\pm$ 0.020± 0.020 0.763 ±0.044$\pm$ 0.044± 0.044 0.829 ±0.012$\pm$ 0.012± 0.012
H4ac 0.372 ±0.016$\pm$ 0.016± 0.016 0.491 ±0.018$\pm$ 0.018± 0.018 0.504 ±0.019$\pm$ 0.019± 0.019 0.525 ±0.240$\pm$ 0.240± 0.240 0.564 ±0.038$\pm$ 0.038± 0.038 0.632 ±0.013$\pm$ 0.013± 0.013
Human TF 0.642 ±0.012$\pm$ 0.012± 0.012 0.633 ±0.015$\pm$ 0.015± 0.015 0.701 ±0.020$\pm$ 0.020± 0.020 - - 0.714 ±0.009$\pm$ 0.009± 0.009
Mouse TF 0.564 ±0.018$\pm$ 0.018± 0.018 0.670 ±0.014$\pm$ 0.014± 0.014 0.680 ±0.015$\pm$ 0.015± 0.015 - - 0.786 ±0.012$\pm$ 0.012± 0.012
Promoter 0.768 ±0.015$\pm$ 0.015± 0.015 0.799 ±0.012$\pm$ 0.012± 0.012 0.774 ±0.019$\pm$ 0.019± 0.019 - - 0.803 ±0.012$\pm$ 0.012± 0.012
Splice Recon. 0.841 ±0.010$\pm$ 0.010± 0.010 0.894 ±0.014$\pm$ 0.014± 0.014 0.850 ±0.020$\pm$ 0.020± 0.020 - - 0.927 ±0.009$\pm$ 0.009± 0.009
Virus Covid 0.555 ±0.017$\pm$ 0.017± 0.017 0.730 ±0.012$\pm$ 0.012± 0.012 0.710 ±0.014$\pm$ 0.014± 0.014 - - 0.706 ±0.015$\pm$ 0.015± 0.015
Overall Avg 0.552 0.636 0.621 0.586 0.610 0.708
We evaluate our models on a comprehensive Genomic Understanding Evaluation (GUE) benchmark comprising tasks from Genomics Benchmark (zhou2023dnabert, ) and Nucleotide Transformer tasks (dalla2024nucleotide, ) by standard LoRA fine-tuning (hu2021lora, ). These include regulatory element classification, histone marker prediction, splice site annotation, and promoter/enhancer recognition tasks across various species. We refer to the best results of baselines in the original paper.
The results are shown in Table 2, comparing models DNABERT, Nucleotide Transformer (NT), DNABERT2, and our proposed TrinityDNA. Overall, TrinityDNA outperforms prior methods across many metrics (e.g., average F1 score, MCC), highlighting our model’s strong ability to capture both local motifs and long-range dependencies in genomic sequences. We observe particularly large improvements in tasks that demand recognition of extended promoter regions or higher-order structural features, aligning with our architectural design for multi-window attention and GRC-based reverse complement awareness. In summary, our results on the GUE benchmark indicate that TrinityDNA not only leverages evolutionary signals during pre-training to handle long sequences effectively but also outperforms existing models in capturing fine-scale regulatory and annotations.
Table 3: Zero-shot performance across DNA, RNA, and protein DMS tasks. All values are performance scores. Bold marks the best value per row, and underlined indicates the second best.
Task Type Task TrinityMicroDNA TrinityDNA EVO EVO2 EVO2 Caduceus
# Params 1B 1B 7B 40B 1B 40M
RNA DMS (Prokaryote) Zhang 0.560 0.476 0.239 0.021 0.152 0.133
Pitt 0.294 0.116 0.021 0.011 0.040 0.175
Hayden 0.365 0.141 0.138 0.065 0.010 0.059
Guy 0.370 0.321 0.214 0.417 0.354 0.019
Kobori 0.569 0.561 0.255 0.226 0.317 0.275
Domingo 0.438 0.372 0.456 0.403 0.315 0.215
Andreasson 0.292 0.276 0.053 0.127 0.036 0.070
Protein DMS (Prokaryote) Firnberg 0.673 0.433 0.552 0.499 0.621 0.018
Jacquier 0.659 0.409 0.471 0.446 0.529 0.023
Kelsic 0.406 0.397 0.321 0.309 0.463 0.047
Weeks 0.573 0.505 0.526 0.492 0.561 0.034
Rockah 0.592 0.411 0.574 0.715 0.657 0.168
Chen 0.383 0.344 0.448 0.630 0.534 0.053
DNA ClinVar 0.629 0.933 0.555 0.950 0.927 0.657
ClinVar-non coding 0.503 0.931 0.397 0.974 0.920 0.601
ClinVar-coding 0.654 0.930 0.484 0.910 0.898 0.699
Protein DMS (Eukaryote) Sun 0.202 0.315 0.314 0.295 0.334 0.097
Tsuboyama 0.595 0.708 0.595 0.773 0.717 0.134
Faure 0.067 0.609 0.284 0.381 0.482 0.041
Average Performance (Prokaryote) 0.475 0.366 0.328 0.335 0.353 0.099
Average Performance (Eukaryote) 0.404 0.699 0.415 0.667 0.670 0.314
4.3.2 Zero-shot Performance
We evaluated our two 1 B-parameter models on 19 zero-shot downstream tasks—covering DNA pathogenicity (CliVar), seven RNA DMS benchmarks and fifteen protein-fitness benchmarks across prokaryotic and eukaryotic organisms—and found that they achieve the best score on 10 of the 19 tasks, outperforming every EVO variant and the 40 M-parameter Caduceus baseline (Table 3). The prokaryote-focused TrinityMicroDNA-1B dominates the prokaryotic regime, winning 8 of 13 prokaryotic tasks and attaining the highest prokaryotic average (0.475) despite its compact size, whereas the multi-species TrinityDNA-1B excels on eukaryotic protein-fitness prediction, delivering the top Faure score and the highest eukaryotic average (0.699), even surpassing the 40 B-parameter EVO2 model (0.667). TrinityDNA also ties state-of-the-art DNA pathogenicity performance by leading on CliVar-coding and placing a close second overall. These complementary strengths underscore the benefit of evolutionary-stage-aware pre-training, and a UMAP projection of genome-level embeddings for ten representative clades (Appendix C) reveals clear taxonomic clustering, indicating that both models learn rich species-specific signals without task-specific fine-tuning.
Table 4: Results of CDS Annotation Benchmark on the filtered RefSeq test set.
Category Models Exact Match 75% Match
Recall Precision F1 Recall Precision F1
Models TrinityMicroDNA-1B 0.775 0.740 0.754 0.826 0.788 0.803
TrinityMicroDNA-470M 0.743 0.623 0.692 0.803 0.693 0.755
TrinityMicroDNA-6M 0.592 0.333 0.488 0.723 0.445 0.524
Caduceus-40M 0.149 0.148 0.140 0.194 0.189 0.180
Pipelines Prodigal 0.832 0.666 0.725 0.909 0.765 0.829
GENSCAN 0.721 0.681 0.702 0.810 0.774 0.799
Glimmer 0.704 0.663 0.688 0.802 0.760 0.780
4.3.3 CDS Annotation Benchmark
We also introduce a novel CDS Annotation Benchmark, aiming to assess long-sequence inference capabilities, practical utility for gene annotation in real-world genomes.
Data Source
From RefSeq, we collect all prokaryotic reference genomes and parse the GenBank annotation files for gene positions/types. This yields token-level labels indicating whether each token belongs to a coding sequence (CDS) with 20k sequence length and, if so, in which strand/direction it is transcribed. The detailed data statistics are described in Appendix C.1.
Results
The results are shown in Table 4.3.2. While Prodigal shows strong recall performance, TrinityMicroDNA-1B delivers the best precision and F1 scores for exact matches, highlighting the model’s strong generalization capabilities across diverse datasets compared to classical pipelines.
5 Conclusion and Limitations
We present TrinityDNA, a foundational DNA model that integrates biologically inspired modules—Groove Fusion and Gated Reverse Complement—alongside a multi-scale attention mechanism to capture both local sequence signals and global genomic context effectively. Built upon an evolutionary training strategy that transitions gradually from prokaryotic to eukaryotic genomes, TrinityDNA gains strong generalization for diverse genomic prediction tasks. To further demonstrate its versatility, we introduce the CDS Annotation Benchmark, which evaluates coding sequence identification and annotation across organisms, providing a realistic standard for genome-scale annotation.
Limitations.
Although evolutionary training excels at long-sequence tasks, it can reduce performance on shorter prokaryotic sequences. Moreover, TrinityDNA has largely been validated on discriminative tasks, leaving its generative potential—such as modeling DNA–protein complexes—unexplored. Future work will refine short-sequence performance, investigate generative capabilities, and scale to larger datasets for broader genomic applications.
Appendix A Pretraining Details
A.1 Experimental Setup Details
In this appendix, we offer comprehensive details regarding the experimental setup for our research.
A.1.1 Hardware Configuration
We conduct our experiments on a cluster of 31 host machines. Each machine is equipped with 8 A100 GPUs, providing substantial parallel computing power. The total number of CPU cores available across the cluster is 128, and the total memory amounts to 1007 GB. The operating system used is Ubuntu with the kernel version 5.4.0 - 139 - generic. Additionally, we utilize RDMA (Remote Direct Memory Access) technology to enhance data transfer efficiency. The DCGM (Data Center GPU Manager) drive version is 525.147.05.
A.1.2 Training Framework
Our training framework is built upon the Megatron and DeepSpeed training frameworks, integrating FlashAttention to accelerate attention computation. It supports DeepSpeed-Ulysess for efficient memory management and large-scale model training, while also incorporating Pipeline Parallelism, Model Parallelism, and Data Parallelism (4D). This unique combination allows us to train models with longer context lengths and more complex architectures without compromising performance. Notably, while DeepSpeed-Ulysess currently does not support Pipeline Parallelism, our framework has successfully integrated both DeepSpeed-Ulysess and Pipeline Parallelism alongside Model and Data Parallelism, enabling a highly scalable and efficient 4D parallelism approach for state-of-the-art model training. For our experiments, we build upon the PyTorch and PyTorch Lightning frameworks, which are widely recognized for their flexibility and efficiency in deep learning research. Our model adopts a Transformer encoder-only architecture. We incorporate several advanced techniques within this architecture:
Activation Function
: We use the GEGLU (Gated Gaussian Error Linear Unit) activation function, which has shown superior performance in handling complex data patterns compared to traditional activation functions.
Normalization Layers
: To stabilize the training process, we employ DeepNorm and LayerNorm. DeepNorm helps in mitigating the vanishing and exploding gradient problems, especially in deep neural networks, while LayerNorm normalizes the input across the feature dimension.
Positional Encoding
: We use RoPE (Rotary Positional Embedding) extended with Dynamic NTK scaling. RoPE is a powerful positional encoding method that can effectively capture the relative position information in sequences, and the Dynamic NTK scaling further enhances its ability to handle long-range dependencies.
Attention Mechanism
: We leverage Flash Attention 2, which significantly reduces the computational complexity of the attention mechanism, enabling faster and more memory-efficient training, especially for long sequences.
A.1.3 Training Optimization
We utilize DeepSpeed - Ulysses for training Transformer models with extremely long sequences. This framework is further modified and optimized based on the pipe mode.
Regarding the training settings, we set the ZERO_STAGE to 1, which helps in reducing memory usage during training by partitioning the optimizer states. We use the BF16 (Brain Floating Point 16) data type for the model parameters to speed up the training process, while the gradient accumulation data type is set to FP32 to maintain numerical stability.
Unless otherwise specified, we use cross-entropy loss as our objective function for training, which is a common choice for classification-related tasks in genomic sequence modeling.
Table 5: Model Configuration of Pretraining
Model Size Sequence Length Layers Hidden Size FFN Hidden Size Num Heads Learning Rate
Baselines 6M 8192 8 256 682 8 1.00E-03
TrinityDNA 6M 8192 8 256 682 8 1.00E-03
Baselines 40M 8192 10 576 1536 8 6.00E-04
TrinityDNA 40M 8192 10 576 1536 8 6.00E-04
Baselines 85M 8192 12 768 2048 12 5.50E-04
TrinityDNA 85M 8192 12 768 2048 12 5.50E-04
Baselines 170M 8192 24 768 2048 16 5.00E-04
TrinityDNA 170M 8192 24 768 2048 16 5.00E-04
Baselines 470M 8192 24 1280 3413 20 4.00E-04
TrinityDNA 470M 8192 24 1280 3413 20 4.00E-04
Baselines 1B 8192 24 2048 5461 32 3.00E-04
TrinityDNA 1B 8192 24 2048 5461 32 3.00E-04
TrinityDNA 1B 30720 24 2048 5461 32 2.00E-04
TrinityDNA 1B 102400 24 2048 5461 32 1.00E-04
A.2 Data
Data Sources
We integrate raw DNA sequence data from the GTDB (Genome Taxonomy Database), IMG (Integrated Microbial Genomes), and RefSeq databases. The OpenGenome dataset was composed by sampling bacterial and archaeal genomes from the GTDB v214.1, curated prokaryotic viruses from IMG/VR v4, and plasmid sequences from IMG/PR, retaining representative genomes for each species. The Multispecies dataset comprises diverse species selected from RefSeq at the genus level, totaling 850 species and about 174 billion nucleotides, ensuring a broad coverage of evolutionary lineages. This data encompasses a wide range of genomic information, including genomic sequences and transcriptomic data. By aggregating data from multiple authoritative sources, we ensure the diversity and comprehensiveness of our dataset, enabling the model to learn a broad spectrum of DNA sequence patterns and biological features.
Data Tokenization
To facilitate model processing, we encode the input DNA sequences using token embedding vectors. Our vocabulary size is set to 5, consisting of the four nucleotide bases A, T, C, and G, and the special character N, which represents an unknown base. This encoding scheme transforms the biological sequences into a numerical format that can be readily processed by our model.
Pretraining Data Preparation
During the pretraining phase, we adopt a masked language model (MLM) strategy. Specifically, we randomly select 15% of the tokens in the DNA sequences as masked tokens. These selected tokens are then replaced with a special $<$mask$>$ token. This process allows the model to learn to predict the original nucleotides based on the context provided by the surrounding unmasked tokens.
However, since the $<$mask$>$ token is typically not present in downstream fine-tuning tasks, we introduce a more sophisticated replacement rule to mitigate the inconsistency between pretraining and fine-tuning. Among the 15% of selected tokens: 1. With an 80% probability, a token is replaced with the $<$mask$>$ token. To further align the long-sequence training process with downstream fine-tuning, there is a 0.02 probability that the replacement ratio can range from 0 to 80%. 2. With a 10% probability, a token is replaced with a randomly chosen token from the vocabulary. 3. With a 10% probability, a token remains unchanged.
Sampling Strategy
Inspiration for our sampling strategy is drawn from related works. Similar to using a single human reference genome and specific training and validation intervals in some studies, our approach is tailored to our integrated dataset. During training, we sample intervals from the integrated sequences and adjust the intervals at both ends to obtain sequences of length L. For the test set, we carefully select specific genomic regions to ensure the reliability of model evaluation. Although we do not follow the exact chromosome selection (chromosomes 14 and X) as in some references, we adopt a similar principle of using non-overlapping sequences of length L to evaluate the model’s performance on unseen data. This sampling strategy helps to ensure that the model is trained on diverse and representative data and can generalize well to new sequences.
A.3 Models
Table 5 presents the model configurations for the pretraining phase, comparing the settings of baseline models and our proposed TrinityDNA model.
Overall Configuration
The table includes models with different parameter scales, denoted as 6M, 40M, 85M, 170M, 470M, and 1B. These models share several common structural features, such as the number of layers, hidden size, feed-forward network (FFN) hidden size, and the number of attention heads. The sequence length for most of the models is set to 8192, except for some TrinityDNA models, where the sequence lengths are 30720 and 102400.
Baseline Models
For the baseline models, as the model scale increases (from 6M to 1B), the number of layers, hidden size, FFN hidden size, and the number of attention heads generally increase. Correspondingly, the learning rate decreases, which is a common practice in deep learning to reset the training process for larger models.
TrinityDNA Models
The TrinityDNA models are designed to have the same basic configurations as the baselines for a fair comparison. For each model scale (e.g., 6M, 40B), the TrinityDNA model has identical hyperparameters to its corresponding baseline model in terms of sequence length, number of layers, hidden size, FFN hidden size, number of attention heads, and learning rate when the sequence length is 8192.
However, when considering the longer sequence lengths of 30720 and 102400 for the 1B TrinityDNA model, the learning rate is further reduced. This adjustment is likely to ensure stable training when dealing with much longer sequences, as longer sequences can introduce more complex dependencies and potentially lead to training instability.
Appendix B Downstream Task Details
In our study, apart from the Zero-shot benchmark, all other downstream tasks are carried out in the form of LoRA (Low-rank Adaptation) fine-tuning [12]. Table 6 presents the detailed LoRA fine-tuning parameters for a variety of tasks, which is essential for understanding the specific configurations and experimental setups of each task.
B.1 LoRA Fine-tuning Parameters on GUE Benchmark [39]
Table 6: LoRA Fine-tuning Parameters
Task lr LoRA rank LoRA alpha batch size
tf_0 0.001 4 8 16
tf_1 0.0001 4 8 16
tf_2 0.001 4 8 16
tf_3 0.0001 4 8 16
tf_4 0.001 4 8 16
mouse_0 0.001 4 8 16
mouse_1 0.0001 24 48 16
mouse_2 0.0001 4 8 16
mouse_3 0.001 4 8 16
mouse_4 0.0001 4 8 16
core_all 0.0001 4 8 16
core_notata 0.0001 4 8 16
core_tata 0.0001 24 48 16
300_all 0.0001 24 48 16
300_notata 0.0001 8 16 16
300_tata 0.001 4 8 16
splice_reconstructed 0.0001 24 48 16
virus_covid 0.0001 4 8 16
H3 0.001 4 8 16
H3K14ac 0.001 4 8 64
H3K36me3 0.0001 96 192 16
H3K4me1 0.001 4 8 64
H3K4me2 0.001 4 8 64
H3K4me3 0.0001 48 96 16
H3K79me3 0.0005 24 48 64
H3K9ac 0.001 4 8 32
H4 0.0001 4 8 32
H4ac 0.001 4 8 32
B.1.1 Task Descriptions
Promoter detection (Human).
This task focuses on detecting proximal promoter regions, the critical genomic sequences that initiate transcription. Because these segments host numerous key regulatory elements, precise identification is essential for advancing our understanding of gene regulatory mechanisms and recognizing the genomic basis of various diseases. Following [7], promoter sequences are split into TATA and non-TATA based on whether they contain a TATA box motif. For each subgroup, we extract the region from 249-249- 249 to +50+50+ 50 base pairs around the transcription start site (TSS) to form the positive (promoter) class. For the negative (non-promoter) class, we randomly select equally sized sequences that (1) contain a TATA motif but lie outside promoter regions (TATA non-promoters), or (2) are formed by random substitutions (non-TATA non-promoters). We also merge both TATA and non-TATA subsets into a combined dataset labeled all.
Core promoter detection (Human).
Similar to the proximal promoter task, this variant targets an even smaller window -34 to +35 base pairs around the TSS to capture the core promoter region. Predicting the core region is more challenging due to the limited context, as it concentrates on the immediate surroundings of the TSS and the start codon.
Transcription factor binding site prediction (Human).
This task involves forecasting transcription factor (TF) binding sites in the human genome. TF binding sites are central to gene expression regulation; accurate prediction helps decode intricate gene regulatory networks and identify therapeutic targets. We use 690 ENCODE ChIP-seq experiments [29], covering 161 TF binding profiles in 91 cell lines. A 101-bp window around the center of each peak defines the positive (TFBS) class, while segments of the same length and matched GC content form the negative (non-TFBS) class. To avoid trivial or overly difficult tasks (e.g., F1 scores $>$ 0.95 or $<$ 0.50), we filter out such datasets and randomly pick 5 from the remaining pool.
Splice site prediction (Human).
This task locates splice donor and acceptor sites in the human genome. Correct splice site identification is crucial for understanding protein diversity and the pathological effects of aberrant splicing. We use a dataset [33] of 400-bp sequences from the Ensembl GRCh38 reference genome. Following [13], we note that models often attain near-perfect results on the original set (10,000 splice donors, acceptors, and non-splice sites), which does not reflect the real difficulty of detecting non-canonical sites. To address this, we iteratively enrich the dataset with adversarial examples (previously unseen false positives) in a hold-out set, making the task substantially harder.
Enhancer–promoter interaction (Human).
This binary classification task aims to identify whether an enhancer interacts with a promoter, a crucial relationship in the human genome that modulates gene expression. The input comprises sequence pairs (enhancer and promoter), and the output is a binary prediction indicating an interaction or lack thereof.
Species classification (Virus & Fungi).
Here, the goal is to classify species based on genomic segments. We build these datasets from viral and fungal reference genomes obtained from GenBank [2].
Transcription factor binding site prediction (Mouse).
Analogous to the human TFBS task, this variant deals with mouse genomes using Mouse ENCODE ChIP-seq data [27]. We generate negative sequences by applying dinucleotide shuffling while preserving frequency. All other settings remain consistent with the human TFBS dataset. As with the human version, we filter out tasks with extreme F1 scores and randomly sample 5 datasets from the remaining collection.
Epigenetic marks prediction (Yeast).
This task forecasts epigenetic modifications (e.g., DNA or histone modifications) in yeast. Such marks affect gene regulation without altering the DNA sequence. We downloaded 10 datasets from [30] and split each into training, validation, and test sets with an 8:1:1 ratio.
Covid variant prediction (Virus).
Focusing on SARS-CoV-2, this task predicts the virus’s variant type from 1000-bp genomic fragments. We gather data from EpiCoV [14] provided by GISAID, covering 9 variants: Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda, and Zeta.
Configurations.
As shown in Tab 6, LoRA Rank and LoRA Alpha LoRA is a method used to fine-tune large pre-trained models more efficiently. The LoRA rank defines the rank of the low-rank matrices employed in the adaptation process, and LoRA alpha serves as a scaling factor. These two parameters jointly control the complexity and expressiveness of the LoRA adaptation. In the table, the LoRA rank ranges from 4 to 96, and the LoRA alpha ranges from 8 to 192. Different tasks use distinct combinations of LoRA rank and LoRA alpha, suggesting that the optimal LoRA configuration depends on the characteristics of each task.
Appendix C Zero-shot Details
C.0.1 Datasets.
DNA pathogenicity (ClinVar).
We follow the pipeline of [21]: single-nucleotide polymorphisms (SNPs) labelled as pathogenic or benign are taken from the ClinVar release, and up to 4k nucleotide flanking windows are retrieved from the GRCh38 human reference genome to form input sequences.
RNA deep-mutational scanning (DMS).
Seven non-coding RNA DMS benchmarks—Zhang, Pitt, Hayden, Guy, Kobori, Domingo and Andreasson—are collected exactly as in [21]. They cover diverse prokaryotic RNAs.
Protein DMS.
Starting from the proteinGYM suite [23], we keep every experiment for which the wild-type codon context can be unambiguously reconstructed. This yields nine prokaryotic sets (Firnberg, Jacquier, Kelsic, Weeks, Rockah, Chen) and one human set (Sun). To broaden the eukaryotic coverage, we further add two recently released human protein scans. For cases where the original authors reported only amino-acid sequences, we rebuild nucleic-acid inputs by selecting the most frequent codon for each amino-acid according to a codon-usage table computed from all CDS regions of GRCh38.
Figure 7: Zero-shot Embedding visualization of different species.
C.0.2 Zero-shot Metric Computation.
For MLM-based models, we employ two metrics: (1) Mean PPL over masked positions: we sequentially mask token(s) and measure the resulting perplexity. (2) Masked Mutation Impact: We compare the PPL difference between wild-type (wt) and mutant (mt) sequences, focusing specifically on masked variant positions. For generative models such as EVO, we rely on likelihood-based scores under the same masking scheme, assessing how well the model anticipates mutations.
Figure 8: Comparison diagrams of two examples in the CDS annotation task. The orange parts are the annotation results of the prodigal (serving as the ground truth), and the red parts are the prediction results of TrinityDNA.
C.1 CDS Benchmark
C.1.1 Data and Configuration
Figure 9: Distribution of taxonomic groups in the training and IID test sets. Our benchmark datasets are constructed from a comprehensive collection of bacterial and archaeal genomes (17,339 genomes across 61 phyla). From these, we selected 35 phyla for training, and the IID test set is randomly sampled from these training phyla, maintaining the same taxonomic distribution. This design enables evaluation of the model’s ability to generalize within known taxonomic groups.
Figure 10: Distribution of taxonomic groups in the OOD test sets. To evaluate model generalization beyond the training distribution, we constructed two distinct OOD test sets: A phylogenetically diverse set of 45 genomes carefully selected from the remaining phyla not used in training, ensuring maximum coverage of untrained phyla.
Data Source.
Obtain all prokaryotic reference genomes from RefSeq, and construct token-level annotation labels based on gene localization and type information in GenBank [2] annotation files.
Data Splits.
RefSeq Dataset: We exclude the 35 training phyla, leaving 26 additional phyla in RefSeq. We randomly sample two species per phylum (or one if only one species exists), yielding 45 genomes total. These serve as an out-of-distribution test set. This benchmark tests both the model’s ability to handle very long sequences and its competence in gene structure recognition. By comparing performance on IID, metagenomic, and cross-phyla sets, we gauge the model’s robustness and generalization in realistic microbial genome annotation scenarios. The training and test data statistics are shown in Figure 9 and 10.
LoRA Tuning Setup
[…]
C.1.2 Evaluation Metrics on the CDS Benchmark
We evaluate gene annotations on the CDS Benchmark using standard metrics of recall, precision, and F1, where
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}},\quad\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}},\quad F_{1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}.$
Here, TP, FP, and FN represent true positives, false positives, and false negatives, respectively.
Exact Match.
A predicted coding sequence (CDS) is counted as a TP under Exact Match if the predicted region’s start and end coordinates exactly match those of the reference annotation. In other words, the tool must recover the full CDS boundaries without offset.
75% Match.
In line with [20], we introduce a less stringent matching criterion. Under 75% Match, a predicted CDS is counted as a TP if:
  1. 1.
    The predicted CDS lies fully within the boundaries of the true (annotated) CDS.
  2. 2.
    The predicted CDS length is at least 75% of the length of the true CDS.
  3. 3.
    The predicted direction is matched.
This relaxes the exact requirement on start/end positions and instead focuses on substantially overlapping the annotated region.
Appendix D Related Work
Long-Sequence Model
Transformers, with their multi-head attention, can capture such dependencies but suffer from quadratic computational complexity with sequence length, limiting their use in long-sequence tasks like genomic analysis [6]. To address this, models like BigBird use sparse attention to expand context size [36]. Based on SSM-derived operators, a line of long-sequence models is proposed to handle long DNA sequences (up to 1M bps) but is unidirectional and not robust to RC inputs [17, 24, 18, 21].
DNA Foundation Model
Early DNA language models such as DNABERT were restricted by Transformer’s complexity, with limited context sizes for DNA sequences [13, 39, 3, 22, 15]. Recent research focuses on integrating biological features. Caduceus (MambaDNA) exploits DNA’s reverse-complement symmetry for better gene regulation modeling [19]. Our TrinityDNA model innovatively adds the Groove Fusion module to capture DNA structural features. In training, traditional models on single-type genomic data have limited generalization. Our evolutionary training strategy, with staged training on different genomes and increasing context lengths, enhances model adaptability and performance.
Impact Statements
The work presented in this paper introduces TrinityDNA, a novel deep-learning model designed to address key challenges in DNA sequence modeling. By incorporating biologically informed components such as Groove Fusion and Gated Reverse Complement (GRC), along with a multi-scale attention mechanism and evolutionary training strategy, TrinityDNA makes significant strides in improving the accuracy and efficiency of genomic sequence analysis.
The impact of TrinityDNA extends beyond academic research in computational genomics. The model’s ability to efficiently capture long-range dependencies in DNA sequences enables more accurate predictions of gene functions, regulatory mechanisms, and other biological insights. This can have profound implications for various fields, including personalized medicine, where understanding genetic variations is crucial for disease prevention and treatment. In addition, by improving the understanding of complex biological systems, TrinityDNA could accelerate advancements in biotechnology and drug development, where genomic data plays a pivotal role in identifying novel therapeutic targets and biomarkers.
Furthermore, the model’s scalability and integration of multi-species genomic data position it as a tool with the potential to advance our understanding of the genomic underpinnings of diverse organisms. This has wide-reaching implications for evolutionary biology, conservation efforts, and the study of microbiomes, where large-scale comparative genomic analysis is essential. By bridging the gap between computational methods and biological insights, TrinityDNA serves as a significant step toward more robust, data-driven approaches to understanding the complexities of life at the genomic level.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Abstract and introduction clearly state the claims made, including the contributions made in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We conclude the limitation part in our conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Section method for a detailed description of our proposed method and implementation details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide relative information in the supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See implementation details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We reported the mean and standard deviation over multiple experimental runs in GUE benchmarks.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe in the experiment section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This work follows the code of ethics of NeurIPS.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Section broader impact in appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited previous works with available codes.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No crowdsourcing experiments in this paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: The participants gave their written consent and were conpensated.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM is only used for correcting the grammar.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 1 Introduction
-The rapid advancements in large-scale, long-sequence modeling, particularly in the realm of Natural Language Processing (NLP) (team2023gemini, ; achiam2023gpt, ), have radically transformed the way we approach complex data.
-Deep learning models, such as Transformers (vaswani2017attention, ), have achieved unprecedented success in tasks that span from language translation to text generation, revolutionizing not just NLP but a variety of other fields.
+The rapid advancements in large-scale, long-sequence modeling, particularly in the realm of Natural Language Processing (NLP) (Team et al.
+2023; Achiam et al.
+2023), have radically transformed the way we approach complex data.
+Deep learning models, such as Transformers (Vaswani et al.
+2017), have achieved unprecedented success in tasks that span from language translation to text generation, revolutionizing not just NLP but a variety of other fields.
 These models have proven their capability to capture intricate dependencies in data, providing solutions to challenges that were once considered insurmountable.
-With these breakthroughs in NLP, there has emerged an exciting opportunity to extend the power of sequence modeling to a completely different domain—genomics—where data shares some key similarities, such as its sequential nature (zhou2015predicting, ).
+With these breakthroughs in NLP, there has emerged an exciting opportunity to extend the power of sequence modeling to a completely different domain—genomics—where data shares some key similarities, such as its sequential nature.
 Genomic data, particularly DNA sequences, consists of extraordinarily long strings of information that encode the fundamental building blocks of life.
-Unlike the highly dense and structured data typically encountered in NLP, genomic sequences are sparse in nature, containing vast stretches of repetition and variability (liu2024genbench, ).
+Unlike the highly dense and structured data typically encountered in NLP, genomic sequences are sparse in nature, containing vast stretches of repetition and variability (Liu et al.
+2024a).
 Despite this, they hold a rich repository of biological information that is crucial for understanding gene functions, regulatory mechanisms, and cellular processes.
@@
 (Right) Radar chart illustrating the state-of-the-art performance on the zero-shot performance of our models versus popular models such as EVO and Caduceus.
-The TrinityMicroDNA model refers to a model trained only on prokaryotic data, which typically performs better on prokaryotic tasks, while TrinityDNA is better on eukaryotic tasks.
 While the parallels between NLP and genomics are evident, directly applying traditional NLP models to genomic sequences proves difficult.
-The sparse, low-density nature of DNA sequences means that existing models often struggle to identify long-range dependencies and interpret the underlying biological structures(mallet2021reverse, ; zhou2021towards, ).
+The sparse, low-density nature of DNA sequences means that existing models often struggle to identify long-range dependencies and interpret the underlying biological structures(Mallet and Vert 2021; Zhou et al.
+2021).
 Moreover, the lack of biologically informed features in current models limits their effectiveness in genomic contexts.
@@
   • New Benchmark for Long-Sequence Inference: We introduce a novel CDS Annotation Benchmark that focuses on gene-structure labeling in prokaryotic genomes, assessing both long-sequence modeling and practical annotation performance.
-In summary, TrinityDNA represents a significant advancement in DNA sequence modeling, addressing critical limitations of existing models through innovative architectural designs, incorporation of biological insights, and robust training strategies.
-By bridging the gap between NLP advancements and genomic data intricacies, TrinityDNA paves the way for more accurate predictions of gene functions, enhanced understanding of genetic regulation mechanisms, and broader applications in genomics research.
 2 Background
-2.1 DNA Terminology for Structures.
+DNA Terminology for Structures.
 Basic Composition
@@
 A DNA sequence can be represented as a string $S=(s_{1},s_{2},\ldots,s_{N})$, where $s_{i}\in\{A,T,C,G\}$ and $N$ denotes the sequence length.
-Genomic sequences are characterized by their low-density encoding and considerable length, often extending to millions of nucleotides, which presents unique challenges for sequence modeling.
 Minor and Major Grooves.
@@
 While less accessible than the major groove, its length is three to five nucleotides.
-Understanding the geometric properties of these grooves is essential for predicting functional regions within the genome.
 Reverse Complement Strands.
 DNA molecules consist of two complementary strands running in opposite directions, a feature known as antiparallel orientation.
-For a DNA sequence $S=(s_{1},s_{2},\ldots,s_{N})$, its reverse complement $S^{R}$ is defined as:
+For a DNA sequence $S=(s_{1},s_{2},\ldots,s_{N})$, its reverse complement $S^{R}$:
 $S^{R}=(s_{N}^{C},s_{N-1}^{C},\ldots,s_{1}^{C})$
@@
 Incorporating reverse complement information into computational models enhances their ability to capture symmetrical and complementary patterns, thereby improving predictions related to gene annotation and regulatory element identification.
-2.2 DNA Long-Sequence Modeling
+DNA Long-Sequence Modeling
 Structured State Space Models (SSMs).
-A prominent class of models for handling long-range dependencies is based on Structured State Space Models (SSMs) (gu2021efficiently, ; gu2021combining, ; gu2022parameterization, ; gupta2022diagonal, ; smith2022simplified, ; dao2022hungry, ).
+A prominent class of models for handling long-range dependencies is based on Structured State Space Models (SSMs) (Gu et al.
+2022a, 2021, b; Gupta et al.
+2022; Smith et al.
+2022; Dao et al.
+2022).
 These models emerge from discretizing a continuous-time linear system:
@@
 HyenaDNA (EVO).
-HyenaDNA (poli2023hyena, ) is an SSM variant, a decoder-style model capable of handling long genomic sequences (e.g., hundreds of thousands of tokens).
+HyenaDNA (Poli et al.
+2023) is an SSM variant, a decoder-style model capable of handling long genomic sequences (e.g., hundreds of thousands of tokens).
 It leverages the Hyena operator, which replaces traditional attention with a fast convolution-based mechanism.
@@
 Caduceus (MambaDNA).
-MambaDNA (schiff2024caduceus, )—also referred to as Caduceus—builds on the selective SSM approach of Mamba and incorporates reverse-complement symmetry, a core property of DNA sequences.
+MambaDNA (Schiff et al.
+2024)—also referred to as Caduceus—builds on the selective SSM approach of Mamba and incorporates reverse-complement symmetry, a core property of DNA sequences.
 In standard Mamba, the module processes sequences in a single direction.
@@
 (2) RC Equivariance: MambaDNA explicitly enforces RC symmetry by taking a sequence and its RC as inputs to the same SSM-based module.
-See more details in Appendix D.
+See more details in Appendix E.
 3 TrinityDNA: Sequence, Structure, and Strategy for DNA Modeling
-[…]
-3.1 Preliminaries
+Figure 2: Comparison of log influential scores log $|\partial yt/\partial xs|$ versus distance (ts)(t-s) on HG-38 (Nguyen et al. 2023).
+Preliminaries
 Lost in the locality.
-While SSMs theoretically excel at handling long sequences, they inherently exhibit a locality bias (wang2024understand, ).
-This issue is amplified in DNA sequence modeling, where dependencies across vast stretches of genetic material need to be captured for accurate biological interpretation.
+While SSMs theoretically excel at handling long sequences, they inherently exhibit a locality bias (Wang et al.
+2024).
+This issue is exacerbated in DNA sequence modeling, where dependencies across vast genomic regions must be captured to enable accurate biological interpretation.
 Specifically, in genomic data, long-range dependencies often span tens of thousands or even hundreds of thousands of base pairs.
-Existing models like SSMs typically focus on local dependencies due to the computational challenges of processing long sequences.
-Consequently, our empirical results in Figure 2 demonstrate that the SSM-based model (Caduceus) lost their focus as sequence length increased, while the full-attention-based model (DNABERT2) suffered from heavy computation.
-These limitations hamper the ability of modern DNA foundation models to fully capture the complexities inherent in DNA sequences, particularly when dealing with biological functions and regulatory mechanisms that depend on large, interconnected genomic regions.
+Existing models, such as SSMs, typically focus on local dependencies due to computational challenges in processing long sequences.
+Consequently, our empirical results in Figure 2 demonstrate that the SSM-based model (Caduceus) lost their focus as the sequence length increased, while the full-attention-based model (DNABERT2) suffered from heavy computation.
+These limitations hinder modern DNA foundation models’ ability to fully capture the inherent complexities of DNA sequences and their interconnections across genomic regions.
 Figure 3: Average attention entropy of full self-attention models as sequence length increases.
@@
 As sequences grow in training, full self-attention “flattens out”: attention scores converge toward a uniform, high-entropy distribution in Figure 3, so every token is weighted almost equally, and useful signals vanish.
-This oversmoothing hits low-density data—images, DNA-hardest, where informative tokens are sparse and far apart.
+This oversmoothing affects low-density data—images, DNA, and the hardest— where informative tokens are sparse and far apart.
 In shorter windows, the same model still shows specialized heads, e.g., retrieval vs.
-induction (xiao2024duo, ), but that diversity collapses at the kilobase scale.
-These findings motivate a multi-window, multi-head scheme that mixes narrow local windows with broader global ones to curb entropy and retain head specialization.
+induction (Xiao et al.
+2024), but that diversity collapses at the kilobase scale.
+These findings motivate a multi-window, multi-head scheme that combines narrow local windows with broader global ones to reduce entropy and maintain head specialization.
+Detailed settings refers to the Appendix.
 Figure 4: Model Architecture of TrinityDNA: The model integrates DNA sequences and structural features by considering its grooves and reverse complementary sequence with shared parameters.
-3.2 Groove Fusion Module
+Groove Fusion Module
 To account for the minor and major grooves in DNA sequences, we propose a Groove Fusion module that combines convolutional operations of varying kernel sizes.
@@
 The output of each convolution operation is fused to capture the multi-scale contextual information necessary for interpreting the structural variations between the major and minor grooves.
-3.3 Sliding Multi-Window Attention
-To overcome the locality bias and attention oversmoothing, we revisit the design of Multi-Head Attention (MHA), focusing on varying attention window sizes for different heads.
+Sliding Multi-Window Attention
+To mitigate locality bias and attention oversmoothing, we revisit the design of Multi-Head Attention (MHA), focusing on varying attention window sizes across heads.
 In traditional MHA, each head typically attends to the entire sequence using a fixed attention window, which can limit the model’s ability to capture dependencies at different scales.
-In the context of DNA sequence modeling, dependencies are not uniform across the sequence.
-For example, short motifs require local attention, while regulatory regions may span longer stretches of DNA.
 To address this, we introduce a multi-scale attention grouping strategy named SMWA.
-In this approach, we assign different attention heads to focus on different scales of dependencies in the DNA sequence, allowing the model to specialize in either local or global dependencies depending on the scale of the feature being modeled.
+In this approach, we assign different attention heads to capture dependencies at different scales in the DNA sequence, enabling the model to specialize in local or global dependencies, depending on the feature scale.
 Formally, we define the window sizes for each attention head $h$ as $L_{h}\in\mathbb{N}$, with $L_{h}$ representing the length of the attention window for head $h$.
-For head $h$, the attention mechanism is computed using a sliding window of size $L_{h}$ over the sequence, allowing the model to focus on different scales.
-Specifically, the attention calculation is constrained to the sliding window around each position, which captures dependencies within a local region of the sequence.
-The sliding window can be defined as:
-$\text{Attn}_{h}(S_{i})=\text{Softmax}\left(\frac{Q_{h}(i)K_{h}(i+[-L_{h},L_{h}])^{T}}{\sqrt{d_{k}}}\right)V_{h}(i+[-L_{h},L_{h}])$
+For head $h$, the attention mechanism is computed using a sliding window of size $L_{h}$ over the sequence, allowing the model to focus on scales:
+$\text{Attn}_{h}(S_{i})=$ $\text{Softmax}\left(\frac{Q_{h}(i)K_{h}(i+[-L_{h},L_{h}])^{T}}{\sqrt{d_{k}}}\right)$
+$V_{h}(i+[-L_{h},L_{h}])$
 where $Q_{h}\in\mathbb{R}^{N\times d_{k}}$ is the query matrix for head $h$, $K_{h}\in\mathbb{R}^{N\times d_{k}}$ is the key matrix, $V_{h}\in\mathbb{R}^{N\times d_{v}}$ is the value matrix, $N$ is the sequence length, and $d_{k}$ and $d_{v}$ are the dimensionality of the key and value vectors, respectively.
-$i$ is the index of the sequence, and $[-L_{h},L_{h}]$ represents the range of indices for the sliding window around $i$.
+$i$ is the index of the sequence, and $L_{h},L_{h}]$ represents the range of indices for the sliding window around $i$.
 This enables each head to specialize in attending to either short-range or long-range dependencies by adjusting $L_{h}$.
@@
 Thus, shorter sequences can be captured by heads with smaller windows, while longer-range dependencies can be modeled by heads with larger windows, enabling the model to capture the hierarchical nature of DNA sequences better.
-3.4 Gated Reverse Complement
+Gated Reverse Complement
+Figure 5: Scaling Behaviors of Our Proposed Model.
+(Left) Evaluation perplexity (PPL) against total FLOPs across multiple architectures, showing consistent improvements to various baselines.
+(Right) Impact of increasing context length (8k, 30k, 100k) on a eukaryotic dataset, where PPL steadily decreases with longer context windows.
 To leverage the reverse complementarity inherent in DNA sequences, we introduce a novel Gated Reverse Complement (GRC) mechanism.
@@
 The GRC mechanism works by feeding both the forward and reverse complement sequences into a shared SMWA-equipped Transformer network $f_{\theta}$, where the outputs are gated using a linear gating mechanism to combine the two representations effectively.
-The gating at the final layer is given by:
+The final layer is:
 $\text{Output}=\text{GRC}(S,S^{R})=$ $f_{\theta}(S)+\sigma(W_{G}\cdot f_{\theta}(\text{Flip}(S^{R})))$
-where $W_{G}$ are learned weights, $\sigma$ represents the sigmoid or identity function, and the Flip operator means to reverse the sequence as the original order.
-This allows the model to simultaneously learn the representations of both forward and reverse sequences, capitalizing on the symmetry of DNA, which is crucial for many biological processes.
-3.5 Evolutionary Training Strategy
+where $W_{G}$ are learned weights, $\sigma$ represents the identity function, and the Flip operator means to reverse the sequence in the original order.
+This allows the model to learn the representations of both sides simultaneously.
+Evolutionary Training Strategy
 The Evolutionary Training Strategy (ETS) approach leverages a two-stage, evolution-inspired training strategy to progressively address the varying complexities of genomic data.
-In the first stage, the model is trained on prokaryotic genomes, which average around 924 base pairs (xu2006genome, ; nguyen2024sequence, ) and present relatively straightforward regulatory architectures.
+In the first stage, the model is trained on prokaryotic genomes, which are relatively straightforward in terms of their regulatory architectures (Xu and Jin 2006; Nguyen et al.
+2024).
 Through this foundational phase, the model captures essential DNA sequence motifs and organizational patterns.
-Subsequently, the second stage introduces eukaryotic genomes, known for their intron-exon structures and gene lengths that can span tens of kilobases (dalla2024nucleotide, ).
-Alongside this transition, the model’s context window is enlarged from 8K to 100K base pairs, accommodating multiple co-expressed genes and complex regulatory elements.
-Figure 5: Scaling Behaviors of Our Proposed Model.
-(Left) Evaluation perplexity (PPL) against total FLOPs across multiple architectures, showing consistent improvements to various baselines.
-(Right) Impact of increasing context length (8k, 30k, 100k) on a eukaryotic dataset, where PPL steadily decreases with longer context windows.
+Subsequently, the second stage introduces eukaryotic genomes, known for their intron-exon structures and gene lengths that can span tens of kilobases (Dalla-Torre et al.
+2024).
+Alongside this transition, the model’s context window is enlarged from 8K to 100K base pairs, accommodating multiple co-expressed genes and regulatory elements.
 4 Experiments
-4.1 Pre-training
+Pre-training
 Data.
-We adopt masked language modeling (MLM) (devlin2018bert, ) and character-level tokenization for all our DNA models.
-Following our Evolutionary training strategy, we conduct pre-training in two phases: Stage 1 (Short-Sequence Pre-training): We use prokaryotic (bacterial and archaeal) DNA data from the OpenGenome dataset introduced in nguyen2024sequence , training on sequences of length 8k to learn fundamental nucleic acid patterns in shorter contexts.
-Stage 2 (Multi-species Post-training): We then continue pre-training on a multi-species collection, the Multispecies dataset, presented in dalla2024nucleotide .
+We adopt masked language modeling (MLM) (Devlin et al.
+2018) and character-level tokenization for all our DNA models.
+Following our Evolutionary training strategy, we conduct pre-training in two phases: Stage 1 (Short-Sequence Pre-training): We use prokaryotic (bacterial and archaeal) DNA data from the OpenGenome dataset introduced in (Nguyen et al.
+2024), training on sequences of length 8k to learn fundamental nucleic acid patterns in shorter contexts.
+Stage 2 (Multi-species Post-training): We then continue pre-training on a multi-species collection, the Multispecies dataset, presented in (Dalla-Torre et al.
+2024).
 The sequences in this stage can be as long as 100k, spanning archaebacteria, fungi, vertebrate genomes, and more.
@@
 The former is trained solely on prokaryotic data, while the latter builds upon the former by post-training on eukaryotic data.
-A “100k” label indicates an increased context window size of up to 100k.
 By transitioning from bacterial genomes to longer eukaryotic genomes, our method achieves broad coverage of genomic features across scales.
-See more details in Appendix A.
+More details in Appendix A.
 Scaling Laws.
@@
 (Right) We examine the effect of increasing context window sizes on a eukaryotic benchmark, finding a steady drop in perplexity when moving from 8k to 30k, and a further substantial improvement at 100k.
+Components W/O W W/O W
+GRC 2.731 2.599 (-0.132) - -
+GFM 2.599 2.534 (-0.065) - -
+SMWA 2.534 2.544 (+0.010) 64.5 44.5 (-31%)
+Table 1: Effect of components on pre-training perplexity (left side) and computational cost (left side, TFLOPs).
 Figure 6: Comprehensive ablation study and efficiency analysis on TrinityDNA.
-4.2 Ablation Study and Analysis
-Table 1: Effect of GRC, GFM, and SMWA on pre-training perplexity.
-Components W/O W
-GRC 2.731 2.599 (-0.132)
-GFM 2.599 2.534 (-0.065)
-SMWA 2.534 2.544 (+0.010)
-We comprehensively analyze our model by answering three key questions on the 470M scale: (1) Are the proposed modules effective?
-(2) How efficient of TrinityDNA for long-sequence DNA computing?
-(3) Why do we use ETS for DNA modeling?
-(1) Effectiveness.
-Table 1 shows that GFM consistently lowers perplexity by modeling the spatial ‘groove’ features of DNA sequences.
-Likewise, incorporating GRC (which captures reverse-complement patterns) yields a notable drop in PPL, reflecting the importance of complementary-strand information.
-Meanwhile, SMWA enables multi-scale context handling, trading off some computational overhead for competitive perplexity across local and longer-range dependencies.
-SMWA also brings some performance gain due to the low-density perplexity of the nucleotide sequence.
-(2) Efficiency.
-The left panel of Figure 6 contrasts token-throughput as we sweep both sequence length and micro-batch size.
-Across every setting, TrinityDNA sits clearly at the top, retaining more than 80% of its short-sequence throughput even at 64 k tokens.
-This robustness comes from its sliding multi-window attention and optimized fused kernels, which keep memory traffic almost flat as context grows.
-By contrast, DNABERT-2 exhibits the steepest decline: its throughput falls by nearly an order of magnitude once the sequence length exceeds 16 k, reflecting the quadratic cost of its vanilla soft-max attention.
-The remaining baselines lie between these two extremes, showing gradual slowdowns that track their respective attention or RNN bottlenecks.
-(3) Training Strategies.
-The right plot in Figure 6 shows a comprehensive ablation study separating the contributions of dataset size and evolutionary training strategy.
-The table shows perplexity scores across different model configurations.
-Key finding: Models initialized with weights from prokaryotic pre-training and then fine-tuned on combined data show better performance than models trained from scratch on the combined dataset.
-This validates both the importance of large, diverse datasets and our evolutionary training approach.
-4.3 Downstream Tasks
-We employ TrinityMicroDNA and TrinityDNA with 1B parameters as our base model for downstream evaluation, and we use LoRA tuning except for the zero-shot task.
-4.3.1 GUE Benchmark
-Table 2: GUE Benchmark performance comparison across various datasets and tasks.
 Tasks DNABERT NT DNABERT2 Caduceus HyenaDNA TrinityDNA
 # Params 86M 2.5B 117M 40M 1.6M 1B
-H3 0.731 ±0.015$\pm$ 0.015± 0.015 0.788 ±0.012$\pm$ 0.012± 0.012 0.783 ±0.014$\pm$ 0.014± 0.014 0.799 ±0.029$\pm$ 0.029± 0.029 0.779 ±0.037$\pm$ 0.037± 0.037 0.814 ±0.014$\pm$ 0.014± 0.014
-H3K14ac 0.401 ±0.018$\pm$ 0.018± 0.018 0.562 ±0.015$\pm$ 0.015± 0.015 0.526 ±0.019$\pm$ 0.019± 0.019 0.541 ±0.212$\pm$ 0.212± 0.212 0.612 ±0.065$\pm$ 0.065± 0.065 0.694 ±0.016$\pm$ 0.016± 0.016
-H3K36me3 0.473 ±0.017$\pm$ 0.017± 0.017 0.620 ±0.012$\pm$ 0.012± 0.012 0.569 ±0.015$\pm$ 0.015± 0.015 0.609 ±0.109$\pm$ 0.109± 0.109 0.613 ±0.041$\pm$ 0.041± 0.041 0.692 ±0.014$\pm$ 0.014± 0.014
-H3K4me1 0.414 ±0.012$\pm$ 0.012± 0.012 0.553 ±0.011$\pm$ 0.011± 0.011 0.505 ±0.015$\pm$ 0.015± 0.015 0.488 ±0.102$\pm$ 0.102± 0.102 0.512 ±0.024$\pm$ 0.024± 0.024 0.611 ±0.015$\pm$ 0.015± 0.015
-H3K4me2 0.323 ±0.014$\pm$ 0.014± 0.014 0.365 ±0.010$\pm$ 0.010± 0.010 0.311 ±0.013$\pm$ 0.013± 0.013 0.388 ±0.101$\pm$ 0.101± 0.101 0.455 ±0.095$\pm$ 0.095± 0.095 0.480 ±0.013$\pm$ 0.013± 0.013
-H3K4me3 0.278 ±0.015$\pm$ 0.015± 0.015 0.403 ±0.011$\pm$ 0.011± 0.011 0.363 ±0.014$\pm$ 0.014± 0.014 0.440 ±0.202$\pm$ 0.202± 0.202 0.549 ±0.056$\pm$ 0.056± 0.056 0.522 ±0.015$\pm$ 0.015± 0.015
-H3K79me3 0.612 ±0.013$\pm$ 0.013± 0.013 0.647 ±0.010$\pm$ 0.010± 0.010 0.674 ±0.012$\pm$ 0.012± 0.012 0.676 ±0.026$\pm$ 0.026± 0.026 0.672 ±0.048$\pm$ 0.048± 0.048 0.741 ±0.014$\pm$ 0.014± 0.014
-H3K9ac 0.512 ±0.015$\pm$ 0.015± 0.015 0.560 ±0.013$\pm$ 0.013± 0.013 0.556 ±0.017$\pm$ 0.017± 0.017 0.604 ±0.048$\pm$ 0.048± 0.048 0.581 ±0.061$\pm$ 0.061± 0.061 0.662 ±0.014$\pm$ 0.014± 0.014
-H4 0.793 ±0.012$\pm$ 0.012± 0.012 0.817 ±0.015$\pm$ 0.015± 0.015 0.807 ±0.011$\pm$ 0.011± 0.011 0.789 ±0.020$\pm$ 0.020± 0.020 0.763 ±0.044$\pm$ 0.044± 0.044 0.829 ±0.012$\pm$ 0.012± 0.012
-H4ac 0.372 ±0.016$\pm$ 0.016± 0.016 0.491 ±0.018$\pm$ 0.018± 0.018 0.504 ±0.019$\pm$ 0.019± 0.019 0.525 ±0.240$\pm$ 0.240± 0.240 0.564 ±0.038$\pm$ 0.038± 0.038 0.632 ±0.013$\pm$ 0.013± 0.013
-Human TF 0.642 ±0.012$\pm$ 0.012± 0.012 0.633 ±0.015$\pm$ 0.015± 0.015 0.701 ±0.020$\pm$ 0.020± 0.020 - - 0.714 ±0.009$\pm$ 0.009± 0.009
-Mouse TF 0.564 ±0.018$\pm$ 0.018± 0.018 0.670 ±0.014$\pm$ 0.014± 0.014 0.680 ±0.015$\pm$ 0.015± 0.015 - - 0.786 ±0.012$\pm$ 0.012± 0.012
-Promoter 0.768 ±0.015$\pm$ 0.015± 0.015 0.799 ±0.012$\pm$ 0.012± 0.012 0.774 ±0.019$\pm$ 0.019± 0.019 - - 0.803 ±0.012$\pm$ 0.012± 0.012
-Splice Recon. 0.841 ±0.010$\pm$ 0.010± 0.010 0.894 ±0.014$\pm$ 0.014± 0.014 0.850 ±0.020$\pm$ 0.020± 0.020 - - 0.927 ±0.009$\pm$ 0.009± 0.009
-Virus Covid 0.555 ±0.017$\pm$ 0.017± 0.017 0.730 ±0.012$\pm$ 0.012± 0.012 0.710 ±0.014$\pm$ 0.014± 0.014 - - 0.706 ±0.015$\pm$ 0.015± 0.015
+H3 0.731 $\pm 0.015$ 0.788 $\pm 0.012$ 0.783 $\pm 0.014$ 0.799 $\pm 0.029$ 0.779 $\pm 0.037$ 0.814 $\pm 0.014$
+H3K14ac 0.401 $\pm 0.018$ 0.562 $\pm 0.015$ 0.526 $\pm 0.019$ 0.541 $\pm 0.212$ 0.612 $\pm 0.065$ 0.694 $\pm 0.016$
+H3K36me3 0.473 $\pm 0.017$ 0.620 $\pm 0.012$ 0.569 $\pm 0.015$ 0.609 $\pm 0.109$ 0.613 $\pm 0.041$ 0.692 $\pm 0.014$
+H3K4me1 0.414 $\pm 0.012$ 0.553 $\pm 0.011$ 0.505 $\pm 0.015$ 0.488 $\pm 0.102$ 0.512 $\pm 0.024$ 0.611 $\pm 0.015$
+H3K4me2 0.323 $\pm 0.014$ 0.365 $\pm 0.010$ 0.311 $\pm 0.013$ 0.388 $\pm 0.101$ 0.455 $\pm 0.095$ 0.480 $\pm 0.013$
+H3K4me3 0.278 $\pm 0.015$ 0.403 $\pm 0.011$ 0.363 $\pm 0.014$ 0.440 $\pm 0.202$ 0.549 $\pm 0.056$ 0.522 $\pm 0.015$
+H3K79me3 0.612 $\pm 0.013$ 0.647 $\pm 0.010$ 0.674 $\pm 0.012$ 0.676 $\pm 0.026$ 0.672 $\pm 0.048$ 0.741 $\pm 0.014$
+H3K9ac 0.512 $\pm 0.015$ 0.560 $\pm 0.013$ 0.556 $\pm 0.017$ 0.604 $\pm 0.048$ 0.581 $\pm 0.061$ 0.662 $\pm 0.014$
+H4 0.793 $\pm 0.012$ 0.817 $\pm 0.015$ 0.807 $\pm 0.011$ 0.789 $\pm 0.020$ 0.763 $\pm 0.044$ 0.829 $\pm 0.012$
+H4ac 0.372 $\pm 0.016$ 0.491 $\pm 0.018$ 0.504 $\pm 0.019$ 0.525 $\pm 0.240$ 0.564 $\pm 0.038$ 0.632 $\pm 0.013$
+Human TF 0.642 $\pm 0.012$ 0.633 $\pm 0.015$ 0.701 $\pm 0.020$ - - 0.714 $\pm 0.009$
+Mouse TF 0.564 $\pm 0.018$ 0.670 $\pm 0.014$ 0.680 $\pm 0.015$ - - 0.786 $\pm 0.012$
+Promoter 0.768 $\pm 0.015$ 0.799 $\pm 0.012$ 0.774 $\pm 0.019$ - - 0.803 $\pm 0.012$
+Splice Recon. 0.841 $\pm 0.010$ 0.894 $\pm 0.014$ 0.850 $\pm 0.020$ - - 0.927 $\pm 0.009$
+Virus Covid 0.555 $\pm 0.017$ 0.730 $\pm 0.012$ 0.710 $\pm 0.014$ - - 0.706 $\pm 0.015$
 Overall Avg 0.552 0.636 0.621 0.586 0.610 0.708
-We evaluate our models on a comprehensive Genomic Understanding Evaluation (GUE) benchmark comprising tasks from Genomics Benchmark (zhou2023dnabert, ) and Nucleotide Transformer tasks (dalla2024nucleotide, ) by standard LoRA fine-tuning (hu2021lora, ).
-These include regulatory element classification, histone marker prediction, splice site annotation, and promoter/enhancer recognition tasks across various species.
-We refer to the best results of baselines in the original paper.
-The results are shown in Table 2, comparing models DNABERT, Nucleotide Transformer (NT), DNABERT2, and our proposed TrinityDNA.
-Overall, TrinityDNA outperforms prior methods across many metrics (e.g., average F1 score, MCC), highlighting our model’s strong ability to capture both local motifs and long-range dependencies in genomic sequences.
-We observe particularly large improvements in tasks that demand recognition of extended promoter regions or higher-order structural features, aligning with our architectural design for multi-window attention and GRC-based reverse complement awareness.
-In summary, our results on the GUE benchmark indicate that TrinityDNA not only leverages evolutionary signals during pre-training to handle long sequences effectively but also outperforms existing models in capturing fine-scale regulatory and annotations.
-Table 3: Zero-shot performance across DNA, RNA, and protein DMS tasks. All values are performance scores. Bold marks the best value per row, and underlined indicates the second best.
+Table 2: GUE Benchmark performance comparison. Metrics are single/multi-label binary classification (MCC).
 Task Type Task TrinityMicroDNA TrinityDNA EVO EVO2 EVO2 Caduceus
@@
 Average Performance (Eukaryote) 0.404 0.699 0.415 0.667 0.670 0.314
-4.3.2 Zero-shot Performance
-We evaluated our two 1 B-parameter models on 19 zero-shot downstream tasks—covering DNA pathogenicity (CliVar), seven RNA DMS benchmarks and fifteen protein-fitness benchmarks across prokaryotic and eukaryotic organisms—and found that they achieve the best score on 10 of the 19 tasks, outperforming every EVO variant and the 40 M-parameter Caduceus baseline (Table 3).
-The prokaryote-focused TrinityMicroDNA-1B dominates the prokaryotic regime, winning 8 of 13 prokaryotic tasks and attaining the highest prokaryotic average (0.475) despite its compact size, whereas the multi-species TrinityDNA-1B excels on eukaryotic protein-fitness prediction, delivering the top Faure score and the highest eukaryotic average (0.699), even surpassing the 40 B-parameter EVO2 model (0.667).
-TrinityDNA also ties state-of-the-art DNA pathogenicity performance by leading on CliVar-coding and placing a close second overall.
-These complementary strengths underscore the benefit of evolutionary-stage-aware pre-training, and a UMAP projection of genome-level embeddings for ten representative clades (Appendix C) reveals clear taxonomic clustering, indicating that both models learn rich species-specific signals without task-specific fine-tuning.
-Table 4: Results of CDS Annotation Benchmark on the filtered RefSeq test set.
+Table 3: Zero-shot performance across DNA, RNA, and protein DMS tasks. ClinVar tasks are binary classification (AUC); others are RNA/protein fitness regression (Spearman).
 Category Models Exact Match 75% Match
 Recall Precision F1 Recall Precision F1
-Models TrinityMicroDNA-1B 0.775 0.740 0.754 0.826 0.788 0.803
+Pre-trained Models TrinityMicroDNA-1B 0.775 0.740 0.754 0.826 0.788 0.803
 TrinityMicroDNA-470M 0.743 0.623 0.692 0.803 0.693 0.755
@@
 Caduceus-40M 0.149 0.148 0.140 0.194 0.189 0.180
-Pipelines Prodigal 0.832 0.666 0.725 0.909 0.765 0.829
+Classical Pipelines Prodigal 0.832 0.666 0.725 0.909 0.765 0.829
 GENSCAN 0.721 0.681 0.702 0.810 0.774 0.799
 Glimmer 0.704 0.663 0.688 0.802 0.760 0.780
-4.3.3 CDS Annotation Benchmark
+Table 4: Results of CDS Annotation Benchmark on the filtered RefSeq test set.
+Ablation Study and Analysis
+(1) Effectiveness.
+Table 1 shows that GFM consistently lowers perplexity by modeling the spatial ‘groove’ features of DNA sequences.
+Likewise, incorporating GRC (captures reverse-complement patterns) yields a notable drop in PPL, reflecting the importance of complementary-strand information.
+Meanwhile, SMWA enables multi-scale context handling, trading off some computational overhead for competitive perplexity across local and longer-range dependencies.
+Although SMWA incurs some performance loss, it is still evident that it saves resources for long sequence training.
+(2) Efficiency.
+The left panel of Figure 6 contrasts token-throughput as we sweep both sequence length and micro-batch size on 1B scale.
+Across every setting, TrinityDNA remains clearly at the top, retaining more than 80% of its short-sequence throughput even at 64k tokens.
+This robustness stems from its sliding multi-window attention and optimized fused kernels, which maintain memory traffic at a nearly constant level as the context grows.
+(3) Training Strategies.
+The right plot in Figure 6 shows a comprehensive ablation study separating the contributions of dataset size and evolutionary training strategy.
+The table shows perplexity scores across different model configurations.
+Key finding: Models initialized with weights from prokaryotic pre-training and then fine-tuned on combined data show better performance than models trained from scratch on the combined dataset.
+This validates both the importance of large, diverse datasets and our EST.
+Downstream Tasks
+We employ TrinityMicroDNA and TrinityDNA with 1B parameters as our base model for downstream evaluation, and we use LoRA tuning except for the zero-shot task.
+GUE Benchmark
+We evaluate our models on a comprehensive Genomic Understanding Evaluation (GUE) benchmark comprising tasks from Genomics Benchmark (Zhou et al.
+2023) and Nucleotide Transformer tasks (Dalla-Torre et al.
+2024) by standard LoRA fine-tuning (Hu et al.
+2021).
+For baselines, results are used in the original paper.
+Results
+In Table 2, we compare models DNABERT, Nucleotide Transformer (NT), DNABERT2, and TrinityDNA.
+Overall, ours outperforms prior methods across many metrics.
+We observe large improvements in tasks that demand recognition of extended promoter regions or higher-order structural features, aligning with our architectural design for multi-window attention and GRC-based reverse complement awareness.
+Zero-shot Performance
+We evaluated our two 1B models on 19 zero-shot downstream tasks—covering DNA pathogenicity (ClinVar), seven RNA DMS benchmarks and fifteen protein-fitness benchmarks across prokaryotic and eukaryotic (Table 3).
+The prokaryote-focused TrinityMicroDNA-1B dominates the prokaryotic regime, winning 8 of 13 prokaryotic tasks and attaining the highest prokaryotic average (0.475), whereas the multi-species TrinityDNA-1B excels on eukaryotic protein-fitness prediction, delivering the top score and the highest eukaryotic average (0.699), even surpassing the 40B EVO2 model (0.667).
+TrinityDNA also excels in state-of-the-art DNA pathogenicity performance, leading in CliVar-coding and ranking second overall.
+These complementary strengths underscore the benefit of ETS, and a UMAP projection of genome-level embeddings for ten representative clades (Appendix D) reveals clear taxonomic clustering, indicating that both models learn rich species signals without fine-tuning.
+CDS Annotation Benchmark
 We also introduce a novel CDS Annotation Benchmark, aiming to assess long-sequence inference capabilities, practical utility for gene annotation in real-world genomes.
-Data Source
 From RefSeq, we collect all prokaryotic reference genomes and parse the GenBank annotation files for gene positions/types.
 This yields token-level labels indicating whether each token belongs to a coding sequence (CDS) with 20k sequence length and, if so, in which strand/direction it is transcribed.
-The detailed data statistics are described in Appendix C.1.
+The detailed data statistics are described in Appendix D.
 Results
-The results are shown in Table 4.3.2.
+The results are shown in Table 4.
 While Prodigal shows strong recall performance, TrinityMicroDNA-1B delivers the best precision and F1 scores for exact matches, highlighting the model’s strong generalization capabilities across diverse datasets compared to classical pipelines.
-5 Conclusion and Limitations
-We present TrinityDNA, a foundational DNA model that integrates biologically inspired modules—Groove Fusion and Gated Reverse Complement—alongside a multi-scale attention mechanism to capture both local sequence signals and global genomic context effectively.
-Built upon an evolutionary training strategy that transitions gradually from prokaryotic to eukaryotic genomes, TrinityDNA gains strong generalization for diverse genomic prediction tasks.
-To further demonstrate its versatility, we introduce the CDS Annotation Benchmark, which evaluates coding sequence identification and annotation across organisms, providing a realistic standard for genome-scale annotation.
-Limitations.
-Although evolutionary training excels at long-sequence tasks, it can reduce performance on shorter prokaryotic sequences.
-Moreover, TrinityDNA has largely been validated on discriminative tasks, leaving its generative potential—such as modeling DNA–protein complexes—unexplored.
-Future work will refine short-sequence performance, investigate generative capabilities, and scale to larger datasets for broader genomic applications.
-Appendix A Pretraining Details
-A.1 Experimental Setup Details
+5 Conclusion
+We present TrinityDNA, a foundational DNA model that integrates biologically inspired modules—Groove Fusion and Gated Reverse Complement—alongside a multi-scale attention mechanism to capture scales of genomic context effectively, built upon an ETS that transitions from prokaryotic to eukaryotic genomes.
+TrinityDNA gains strong generalization for diverse genomic prediction tasks.
+We also introduce the CDS Annotation Benchmark, which evaluates coding sequence identification across organisms and provides a realistic standard for genome-scale annotation.
+Appendix A Appendix of TrinityDNA
+Appendix B Experimental Setups
 In this appendix, we offer comprehensive details regarding the experimental setup for our research.
-A.1.1 Hardware Configuration
+Hardware Configuration
 We conduct our experiments on a cluster of 31 host machines.
@@
 The DCGM (Data Center GPU Manager) drive version is 525.147.05.
-A.1.2 Training Framework
+Details for Log Influential Score Analysis
+The analysis of log influential scores was designed to quantitatively measure how the influence of an input token on an output token diminishes with increasing distance between them.
+This helps to empirically visualize the locality bias of different model architectures.
+  1. 1
+    Models and Dataset: We utilized the official pretrained checkpoints for the SSM-based model Caduceus-Phage and the full-attention model DNABERT2.
+The analysis was conducted on the human reference genome, HG-38.
+  2. 2
+    Sequence Sampling: We randomly selected a contiguous sequence of 20,000 base pairs from the HG-38 genome to serve as the input for the models.
+  3. 3
+    Influence Calculation: The influential score is defined as the gradient of a model’s output with respect to its input, specifically $|\partial y_{t}/\partial x_{s}|$.
+This value quantifies the extent to which a change in the input token at position $s$ affects the model’s prediction at position $t$.
+    • To compute this, we performed a forward pass with the 20k-token sequence.
+    • We then used backpropagation to calculate the gradients of the output logits ($y_{t}$) with respect to the input embeddings ($x_{s}$) for a chosen target position $t$ near the end of the sequence.
+    • We computed the absolute value of these gradients and took their logarithm, yielding the log influential score, $\log|\partial y_{t}/\partial x_{s}|$.
+  4. 4
+    Plotting: The scores were plotted against the relative distance between the input and output tokens, (ts)(t-s).
+The curve for each model represents the average scores over multiple runs with different randomly sampled sequences.
+The vertical dashed line labeled “OOM at 16k” indicates the point at which DNABERT2 failed to process the sequence due to an out-of-memory error, a known limitation of its quadratic attention mechanism.
+Details for Average Attention Entropy Analysis
+The experiment on average attention entropy was conducted to demonstrate the “oversmoothing” phenomenon in standard Transformer models, where attention distributions become increasingly uniform as sequence lengths grow.
+  1. 1
+    Model Architecture: We constructed a standard two-layer Transformer encoder with full self-attention.
+The model was intentionally kept simple to isolate the effect of sequence length on the attention mechanism itself, without confounding factors from more complex architectures.
+  2. 2
+    Datasets and Training: We trained two separate instances of this model from scratch:
+    • One on a large corpus of English text from Wikipedia (Text Sequence).
+    • Another on genomic data from the HG-38 reference genome (Nucleotide Sequence).
+    This dual-dataset approach allowed us to compare the oversmoothing effect across different data modalities.
+  3. 3
+    Entropy Calculation: After training, the models were evaluated on sequences of varying lengths, from $2^{10}$ (1024) to $2^{14}$ (16384) tokens.
+    • For each input sequence, we extracted the attention probability distributions, $p$, from every attention head at every position.
+    • We then calculated the Shannon entropy for each distribution using the formula $H(p)=-\sum_{i}p_{i}\log_{2}p_{i}$.
+A higher entropy value indicates a more uniform (or ”smoother”) distribution, while a lower entropy value suggests a more focused, spiky distribution.
+    • The final ”Avg. Attn Entropy” value for each sequence length shown on the y-axis is the mean entropy calculated across all attention heads and all token positions in the sequence.
+  4. 4
+    Plotting: The chart plots the average attention entropy as a function of sequence length.
+The x-axis is on a logarithmic scale ($2^{N}$) to clearly show the trend across several orders of magnitude in sequence length.
+Training Framework
 Our training framework is built upon the Megatron and DeepSpeed training frameworks, integrating FlashAttention to accelerate attention computation.
@@
 : We leverage Flash Attention 2, which significantly reduces the computational complexity of the attention mechanism, enabling faster and more memory-efficient training, especially for long sequences.
-A.1.3 Training Optimization
+Training Optimization
 We utilize DeepSpeed - Ulysses for training Transformer models with extremely long sequences. This framework is further modified and optimized based on the pipe mode.
@@
 Table 5: Model Configuration of Pretraining
-Model Size Sequence Length Layers Hidden Size FFN Hidden Size Num Heads Learning Rate
-Baselines 6M 8192 8 256 682 8 1.00E-03
-TrinityDNA 6M 8192 8 256 682 8 1.00E-03
-Baselines 40M 8192 10 576 1536 8 6.00E-04
-TrinityDNA 40M 8192 10 576 1536 8 6.00E-04
-Baselines 85M 8192 12 768 2048 12 5.50E-04
-TrinityDNA 85M 8192 12 768 2048 12 5.50E-04
-Baselines 170M 8192 24 768 2048 16 5.00E-04
-TrinityDNA 170M 8192 24 768 2048 16 5.00E-04
-Baselines 470M 8192 24 1280 3413 20 4.00E-04
-TrinityDNA 470M 8192 24 1280 3413 20 4.00E-04
-Baselines 1B 8192 24 2048 5461 32 3.00E-04
-TrinityDNA 1B 8192 24 2048 5461 32 3.00E-04
-TrinityDNA 1B 30720 24 2048 5461 32 2.00E-04
-TrinityDNA 1B 102400 24 2048 5461 32 1.00E-04
-A.2 Data
+Size Sequence Length Layers Hidden Size FFN Hidden Size Num Heads Learning Rate
+6M 8192 8 256 682 8 1.00E-03
+40M 8192 10 576 1536 8 6.00E-04
+85M 8192 12 768 2048 12 5.50E-04
+170M 8192 24 768 2048 16 5.00E-04
+470M 8192 24 1280 3413 20 4.00E-04
+1B 8192 24 2048 5461 32 3.00E-04
+1B 30720 24 2048 5461 32 2.00E-04
+1B 102400 24 2048 5461 32 1.00E-04
+Data
 Data Sources
-We integrate raw DNA sequence data from the GTDB (Genome Taxonomy Database), IMG (Integrated Microbial Genomes), and RefSeq databases.
-The OpenGenome dataset was composed by sampling bacterial and archaeal genomes from the GTDB v214.1, curated prokaryotic viruses from IMG/VR v4, and plasmid sequences from IMG/PR, retaining representative genomes for each species.
-The Multispecies dataset comprises diverse species selected from RefSeq at the genus level, totaling 850 species and about 174 billion nucleotides, ensuring a broad coverage of evolutionary lineages.
-This data encompasses a wide range of genomic information, including genomic sequences and transcriptomic data.
-By aggregating data from multiple authoritative sources, we ensure the diversity and comprehensiveness of our dataset, enabling the model to learn a broad spectrum of DNA sequence patterns and biological features.
+The OpenGenome dataset was composed by sampling bacterial and archaeal genomes from the GTDB v214.1, curated prokaryotic viruses from IMG/VR v4, and plasmid sequences from IMG/PR data databases.
 Data Tokenization
@@
 Specifically, we randomly select 15% of the tokens in the DNA sequences as masked tokens.
-These selected tokens are then replaced with a special $<$mask$>$ token.
+These selected tokens are then replaced with a special <<mask>> token.
 This process allows the model to learn to predict the original nucleotides based on the context provided by the surrounding unmasked tokens.
-However, since the $<$mask$>$ token is typically not present in downstream fine-tuning tasks, we introduce a more sophisticated replacement rule to mitigate the inconsistency between pretraining and fine-tuning.
+However, since the <<mask>> token is typically not present in downstream fine-tuning tasks, we introduce a more sophisticated replacement rule to mitigate the inconsistency between pretraining and fine-tuning.
 Among the 15% of selected tokens: 1.
-With an 80% probability, a token is replaced with the $<$mask$>$ token.
+With an 80% probability, a token is replaced with the <<mask>> token.
 To further align the long-sequence training process with downstream fine-tuning, there is a 0.02 probability that the replacement ratio can range from 0 to 80%.
@@
 This sampling strategy helps to ensure that the model is trained on diverse and representative data and can generalize well to new sequences.
-A.3 Models
+Models
 Table 5 presents the model configurations for the pretraining phase, comparing the settings of baseline models and our proposed TrinityDNA model.
@@
 This adjustment is likely to ensure stable training when dealing with much longer sequences, as longer sequences can introduce more complex dependencies and potentially lead to training instability.
-Appendix B Downstream Task Details
-In our study, apart from the Zero-shot benchmark, all other downstream tasks are carried out in the form of LoRA (Low-rank Adaptation) fine-tuning [12].
+Appendix C Downstream Task Details
+In our study, apart from the Zero-shot benchmark, all other downstream tasks are carried out in the form of LoRA (Low-rank Adaptation) fine-tuning (Hu et al.
+2021).
 Table 6 presents the detailed LoRA fine-tuning parameters for a variety of tasks, which is essential for understanding the specific configurations and experimental setups of each task.
-B.1 LoRA Fine-tuning Parameters on GUE Benchmark [39]
+LoRA Fine-tuning Parameters on GUE Benchmark (Zhou et al. 2023)
 Table 6: LoRA Fine-tuning Parameters
@@
 H4ac 0.001 4 8 32
-B.1.1 Task Descriptions
+Task Descriptions
 Promoter detection (Human).
@@
 Because these segments host numerous key regulatory elements, precise identification is essential for advancing our understanding of gene regulatory mechanisms and recognizing the genomic basis of various diseases.
-Following [7], promoter sequences are split into TATA and non-TATA based on whether they contain a TATA box motif.
-For each subgroup, we extract the region from 249-249- 249 to +50+50+ 50 base pairs around the transcription start site (TSS) to form the positive (promoter) class.
+Following (Dreos et al.
+2013), promoter sequences are split into TATA and non-TATA based on whether they contain a TATA box motif.
+For each subgroup, we extract the region from 249-249 to +50+50 base pairs around the transcription start site (TSS) to form the positive (promoter) class.
 For the negative (non-promoter) class, we randomly select equally sized sequences that (1) contain a TATA motif but lie outside promoter regions (TATA non-promoters), or (2) are formed by random substitutions (non-TATA non-promoters).
@@
 TF binding sites are central to gene expression regulation; accurate prediction helps decode intricate gene regulatory networks and identify therapeutic targets.
-We use 690 ENCODE ChIP-seq experiments [29], covering 161 TF binding profiles in 91 cell lines.
+We use 690 ENCODE ChIP-seq experiments (The ENCODE Project Consortium 2012), covering 161 TF binding profiles in 91 cell lines.
 A 101-bp window around the center of each peak defines the positive (TFBS) class, while segments of the same length and matched GC content form the negative (non-TFBS) class.
-To avoid trivial or overly difficult tasks (e.g., F1 scores $>$ 0.95 or $<$ 0.50), we filter out such datasets and randomly pick 5 from the remaining pool.
+To avoid trivial or overly difficult tasks (e.g., F1 scores >> 0.95 or << 0.50), we filter out such datasets and randomly pick 5 from the remaining pool.
 Splice site prediction (Human).
@@
 Correct splice site identification is crucial for understanding protein diversity and the pathological effects of aberrant splicing.
-We use a dataset [33] of 400-bp sequences from the Ensembl GRCh38 reference genome.
-Following [13], we note that models often attain near-perfect results on the original set (10,000 splice donors, acceptors, and non-splice sites), which does not reflect the real difficulty of detecting non-canonical sites.
+We use a dataset (Wang et al.
+2019) of 400-bp sequences from the Ensembl GRCh38 reference genome.
+Following (Ji et al.
+2021), we note that models often attain near-perfect results on the original set (10,000 splice donors, acceptors, and non-splice sites), which does not reflect the real difficulty of detecting non-canonical sites.
 To address this, we iteratively enrich the dataset with adversarial examples (previously unseen false positives) in a hold-out set, making the task substantially harder.
@@
 Species classification (Virus & Fungi).
-Here, the goal is to classify species based on genomic segments. We build these datasets from viral and fungal reference genomes obtained from GenBank [2].
+Here, the goal is to classify species based on genomic segments. We build these datasets from viral and fungal reference genomes obtained from GenBank (Benson et al. 2012).
 Transcription factor binding site prediction (Mouse).
-Analogous to the human TFBS task, this variant deals with mouse genomes using Mouse ENCODE ChIP-seq data [27].
+Analogous to the human TFBS task, this variant deals with mouse genomes using Mouse ENCODE ChIP-seq data (Stamatoyannopoulos et al.
+2012).
 We generate negative sequences by applying dinucleotide shuffling while preserving frequency.
@@
 Such marks affect gene regulation without altering the DNA sequence.
-We downloaded 10 datasets from [30] and split each into training, validation, and test sets with an 8:1:1 ratio.
+We downloaded 10 datasets from the following link: http://www.jaist.ac.jp/tran/nucleosome/members.htm and split each dataset into training, validation, and test sets at an 8:1:1 ratio.
 Covid variant prediction (Virus).
 Focusing on SARS-CoV-2, this task predicts the virus’s variant type from 1000-bp genomic fragments.
-We gather data from EpiCoV [14] provided by GISAID, covering 9 variants: Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda, and Zeta.
+We gather data from EpiCoV (Khare et al.
+2021) provided by GISAID, covering 9 variants: Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda, and Zeta.
 Configurations.
@@
 Different tasks use distinct combinations of LoRA rank and LoRA alpha, suggesting that the optimal LoRA configuration depends on the characteristics of each task.
-Appendix C Zero-shot Details
-C.0.1 Datasets.
+Appendix D Zero-shot Details
+Datasets.
 DNA pathogenicity (ClinVar).
-We follow the pipeline of [21]: single-nucleotide polymorphisms (SNPs) labelled as pathogenic or benign are taken from the ClinVar release, and up to 4k nucleotide flanking windows are retrieved from the GRCh38 human reference genome to form input sequences.
+We follow the pipeline of (Nguyen et al.
+2024): single-nucleotide polymorphisms (SNPs) labelled as pathogenic or benign are taken from the ClinVar release, and up to 4k nucleotide flanking windows are retrieved from the GRCh38 human reference genome to form input sequences.
 RNA deep-mutational scanning (DMS).
-Seven non-coding RNA DMS benchmarks—Zhang, Pitt, Hayden, Guy, Kobori, Domingo and Andreasson—are collected exactly as in [21]. They cover diverse prokaryotic RNAs.
+Seven non-coding RNA DMS benchmarks—Zhang, Pitt, Hayden, Guy, Kobori, Domingo and Andreasson—are collected exactly as in (Nguyen et al. 2024). They cover diverse prokaryotic RNAs.
 Protein DMS.
-Starting from the proteinGYM suite [23], we keep every experiment for which the wild-type codon context can be unambiguously reconstructed.
+Starting from the proteinGYM suite (Notin et al.
+2023), we keep every experiment for which the wild-type codon context can be unambiguously reconstructed.
 This yields nine prokaryotic sets (Firnberg, Jacquier, Kelsic, Weeks, Rockah, Chen) and one human set (Sun).
@@
 Figure 7: Zero-shot Embedding visualization of different species.
-C.0.2 Zero-shot Metric Computation.
+Zero-shot Metric Computation.
 For MLM-based models, we employ two metrics: (1) Mean PPL over masked positions: we sequentially mask token(s) and measure the resulting perplexity.
@@
 Figure 8: Comparison diagrams of two examples in the CDS annotation task.
-The orange parts are the annotation results of the prodigal (serving as the ground truth), and the red parts are the prediction results of TrinityDNA.
-C.1 CDS Benchmark
-C.1.1 Data and Configuration
+The orange parts represent the annotation results of Prodigal (serving as the ground truth), and the red parts represent the prediction results of TrinityDNA.
+CDS Benchmark
+Data and Configuration
 Figure 9: Distribution of taxonomic groups in the training and IID test sets.
@@
 Data Source.
-Obtain all prokaryotic reference genomes from RefSeq, and construct token-level annotation labels based on gene localization and type information in GenBank [2] annotation files.
+Obtain all prokaryotic reference genomes from RefSeq, and construct token-level annotation labels based on gene localization and type information in GenBank (Benson et al. 2012) annotation files.
 Data Splits.
@@
 LoRA Tuning Setup
-[…]
-C.1.2 Evaluation Metrics on the CDS Benchmark
+We adopt Low-Rank Adaptation (LoRA) (Hu et al.
+2021) to fine-tune our model on the CDS Benchmark.
+Specifically, we use a learning rate of $r=0.001$ and a batch size of 32.
+For LoRA hyperparameters, we set the low-rank dimension $\text{LoRA-r}=4$ and the scaling factor $\text{LoRA-}\alpha=32$.
+Furthermore, we apply LoRA to two additional MLP layers in the classification head, whose hidden dimensions are 256 and 128, respectively.
+By decoupling the base model parameters from the low-rank adaptation parameters, LoRA allows us to retain the expressive capacity of the underlying large model while conferring adaptability to new domains with minimal additional parameters.
+Evaluation Metrics on the CDS Benchmark
 We evaluate gene annotations on the CDS Benchmark using standard metrics of recall, precision, and F1, where
@@
 75% Match.
-In line with [20], we introduce a less stringent matching criterion. Under 75% Match, a predicted CDS is counted as a TP if:
+In line with (Nebrao et al. 2021), we introduce a less stringent matching criterion. Under 75% Match, a predicted CDS is counted as a TP if:
   1. 1.
@@
 This relaxes the exact requirement on start/end positions and instead focuses on substantially overlapping the annotated region.
-Appendix D Related Work
+Appendix E Related Work
 Long-Sequence Model
-Transformers, with their multi-head attention, can capture such dependencies but suffer from quadratic computational complexity with sequence length, limiting their use in long-sequence tasks like genomic analysis [6].
-To address this, models like BigBird use sparse attention to expand context size [36].
-Based on SSM-derived operators, a line of long-sequence models is proposed to handle long DNA sequences (up to 1M bps) but is unidirectional and not robust to RC inputs [17, 24, 18, 21].
+Transformers, with their multi-head attention, can capture such dependencies but suffer from quadratic computational complexity with sequence length, limiting their use in long-sequence tasks like genomic analysis (Devlin et al.
+2018).
+To address this, models like BigBird use sparse attention to expand context size (Zaheer et al.
+2020).
+Based on SSM-derived operators, a line of long-sequence models is proposed to handle long DNA sequences (up to 1M bps) but is unidirectional and not robust to RC inputs (Liu et al.
+2024b; Poli et al.
+2023; Liu et al.
+2024c; Nguyen et al.
+2024).
 DNA Foundation Model
-Early DNA language models such as DNABERT were restricted by Transformer’s complexity, with limited context sizes for DNA sequences [13, 39, 3, 22, 15].
+Early DNA language models such as DNABERT were restricted by Transformer’s complexity, with limited context sizes for DNA sequences (Ji et al.
+2021; Zhou et al.
+2023; Dalla-Torre et al.
+2023; Nguyen et al.
+2023; Li et al.
+2024).
 Recent research focuses on integrating biological features.
-Caduceus (MambaDNA) exploits DNA’s reverse-complement symmetry for better gene regulation modeling [19].
+Caduceus (MambaDNA) exploits DNA’s reverse-complement symmetry for better gene regulation modeling (Mallet and Vert 2021).
 Our TrinityDNA model innovatively adds the Groove Fusion module to capture DNA structural features.
@@
 By bridging the gap between computational methods and biological insights, TrinityDNA serves as a significant step toward more robust, data-driven approaches to understanding the complexities of life at the genomic level.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Abstract and introduction clearly state the claims made, including the contributions made in the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We conclude the limitation part in our conclusion section.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not include theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: See Section method for a detailed description of our proposed method and implementation details.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We provide relative information in the supplemental material.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: See implementation details.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We reported the mean and standard deviation over multiple experimental runs in GUE benchmarks.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We describe in the experiment section.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: This work follows the code of ethics of NeurIPS.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: See Section broader impact in appendix.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This paper poses no such risks
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We have cited previous works with available codes.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: This paper does not release new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: No crowdsourcing experiments in this paper.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [Yes]
-  74. Justification: The participants gave their written consent and were conpensated.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: LLM is only used for correcting the grammar.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.